### v3

# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

Abstract
This paper is focused on statistical learning from data that come as probability measures. In this setting, popular approaches consist in embedding such data into a Hilbert space with either Linearized Optimal Transport or Kernel Mean Embedding. However, the cost of computing such embeddings prohibits their direct use in large-scale settings. We study two methods based on measure quantization for approximating input probability measures with discrete measures of small-support size. The first one is based on optimal quantization of each input measure, while the second one relies on mean-measure quantization. We study the consistency of such approximations, and its implication for scalable embeddings of probability measures into a Hilbert space at a low computational cost. Finally, we illustrate our methods and compare them with existing approaches through various numerical experiments.
1 Introduction
Machine Learning (ML) techniques that model data as a set of N probability measures play a crucial role in various applied fields, including signal and image processing, computer vision, and computational biology [30, 36, 38, 28]. In particular, this framework includes distribution regression [41, 47, 6, 5, 35] for which the predictors are probability measures and the responses are scalar. Principal Component Analysis (PCA) of probability measures [52, 46, 8, 11] is also a key problem for dimensionality reduction of distributional data [3]. Performing these and other standard ML tasks on probability measures is challenging since most algorithms are designed to deal with $N$ points in a Euclidean space rather than $N$ probability distributions. However, as most machine learning methods for data analysis depend on the notion of inner-product, a common approach is to embed distributional data into a Hilbert space. To achieve this, there are two popular embeddings which stand out in the literature. The first one is the Linearized Optimal Transport (LOT) embedding [52, 17, 37], which arises from the theory of Optimal Transport (OT) [45, 40] and leverages the Riemannian-like geometry of the space of probability measures endowed with the Wasserstein distance [2]. The second one, known as Kernel Mean Embedding (KME) [38], relies on the use of kernel methods to map probability measures into a Reproducing Kernel Hilbert Space (RKHS).
However, the computational and storing costs of these embeddings make them impractical when dealing with probability measures with large support. This is often the case when observing $N$ empirical measures on point clouds $X^{(i)}=(X^{(i)}_{1},\cdots,X^{(i)}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}},1\leq i\leq N$, with a large number $m_{i}$ of observations. Such datasets are frequently found in flow cytometry [34], where observations collected from $N$ patients represent a considerable amount of cells, each characterized by $d$ bio-markers. For these single-cell data, one usually encounters point clouds of thousands to millions of events (that is $m_{i}\geq 10^{5}$) living in a feature space of dimension $d$ larger than $10$. Using either the LOT embedding or the KME from such raw data becomes critical as these approaches suffer from high computational costs as soon as the number $m_{i}$ of points per clouds is larger than a few thousands.
Given a set of $N$ probability measures with large support size, how to efficiently compute an embedding into a Hilbert space that is statistically consistent with the embedding derived directly from the raw data ?
1.1 Main contributions
In this paper, we consider the problem of embbeding a set of $d$-dimensional input probability measures $(\mu^{(i)})_{i=1}^{N}$ into a Hilbert space at a low computational cost. To that end, we propose to employ a preliminary $K$-quantization step that is either based on optimal quantization of each input measure $\mu^{(i)}$ or on the quantization of the mean measure $\bar{\mu}=\frac{1}{N}\sum_{i=1}^{n}\mu^{(i)}$ as used in [20] for single-cell data analysis. The aim of this $K$-quantization step is to approximate $(\mu^{(i)})_{i=1}^{N}$ by discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$ with supports of size $K$, with $K$ typically small. Using the theory of measure quantization [23, 39], we validate both quantization approaches by showing (see Theorem 3.4 below):
$\mathcal{W}_{2}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{\mu^{(i)}},\frac{1}{N}\sum_{i=1}^{N}\delta_{\nu^{(i)}_{K}}\right)=\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right)\mbox{ as }K\to+\infty,$ (1.1)
where $\mathcal{W}_{2}$ denotes the $2$-Wasserstein distance (2.3) on $\mathcal{P}(\mathcal{P}(\mathcal{X}))$, the set of probability distributions over $\mathcal{P}(\mathcal{X})$, which is itself the set of probability measures with support included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$.
The asymptotic result (1.1) allows to show the convergence of numerous statistics computed from the $\nu^{(i)}_{K}$’s to corresponding quantities for the $\mu^{(i)}$’s as $K\to+\infty$. These include the Wasserstein barycenter and measures of statistical dispersion, from which clustering result can also be derived. In addition, we establish consistency results for quantities computed from the embeddings of the discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$ into a Hilbert space as $K$ increases, using either the LOT or KME framework. Precisely, we study the Gram matrix of pairwise inner-products of the embedded measures, which is a standard quantity used in machine learning applications. Within the LOT framework, we further prove the consistency of PCA computed from the embedded measures, corresponding to log-PCA [19] in the Wasserstein space.
Finally, the soundness and consistency of our method is illustrated with numerical experiments on synthetic and real datasets. We also show that the method based on mean-measure quantization has computational advantages over optimal quantization of each input measure while preserving satisfactory performances, which justifies its use in large scale settings. Finally, we compare our methods to classical embedding techniques for probability measures.
1.2 Related works
While quantization allows to approximate probability measures with a small set of points, other methods also aim at summarizing a dataset with representative samples. Within the framework of coresets [27, 15], one selects a subset of points such that solving a particular problem on this subset yields similar results than solving the problem on the entire dataset. In order to reduce the computational complexity of Gaussian Processes (GP) model, the principle of inducing variables, see e.g. [49], also allows for an approximation of the posterior by choosing a set of representative points and conditioning the GP on these points.
In [14, 44], quantization is employed to embed a set of $N$ probability measures into a finite-dimensional Euclidean space through measure vectorization. More precisely, given $N$ input measures $\mu^{(i)}$, a quantization of the mean measure $\bar{\mu}=\frac{1}{N}\sum_{i=1}^{n}\mu^{(i)}$ by $K$ centers $x_{1},\ldots,x_{K}$ in $\mathbb{R}^{d}$ is first done. Then, they map each measure $\mu^{(i)}$ to $v^{(i)}=(v^{(i)}_{1},\cdots,v^{(i)}_{K})$ a vector of the convex space $\mathbb{R}_{+}^{K}$, where $v^{(i)}_{k}$ roughly represents the mass of the the measure $\mu^{(i)}$ distributed around the center $x_{k}$. Yet, this embedding does not take into account the relative positions of the $K$ centers, and consistency in the sense (1.1) is not shown as we propose in this paper by endowing the set of quantized measures $\nu^{(i)}_{K}$ with the Wasserstein distance (2.2).
In [12], the authors tackle the problem of computing the KME of a probability distribution $\mu$ for which $m$ samples $X_{1},\ldots,X_{m}$ are available. They introduce an estimator of the KME of $\mu$ based on Nyström approximation that can be computed efficiently using a small random subset from the data. Their theoretical and empirical results show that this approach yields a consistent estimator of the maximum mean discrepancy distance between the KMEs of $\mu$ and $\hat{\mu}_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{j}}$ at a low computational cost. However, this Nyström approximation has not been studied for constructing a consistent LOT embedding estimator.
Finally, the benefits of a preliminary quantization step have been studied in [7] to improve the standard plug-in estimator of the OT cost between two probability measures. Still, the simultaneous quantization of $N$ probability measures for the purpose of constructing consistent and scalable embeddings has not been considered so far.
1.3 Organization of the paper
Section 2 is devoted to some background on OT, the LOT and KME embeddings and the quantization principle. In Section 3, we describe our two quantization methods of a set of $N$ probability measures, and we prove their convergence in the sense of (1.1). In Section 4, we apply this result to derive the consistency of statistics derived from the quantized measures. In particular, we prove the consistency of various machine learning methods that take as inputs either the LOT embedding or KME of the discrete measures $(\nu_{K}^{(i)})_{i=1}^{N}$. Section 5 reports the results of numerical experiments using synthetic and real data, and the computational cost of both methods are discussed and compared. The paper ends with a conclusion in Section 6. All proofs are deferred to two technical Appendices A and B, and additional numerical experiments are given in Appendix C.
2 Background
Optimal transport.
Let $\rho$ and $\mu$ be two probability measures with support included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$. For the quadratic cost, the OT problem between $\rho$ and $\mu$ is:
$\min\limits_{\pi\in\Pi(\rho,\mu)}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi(x,y),$ (2.1)
where $\Pi(\rho,\mu)$ is the set of probability measures (or transport plans) on $\mathcal{X}\times\mathcal{X}$ with marginals $\rho$ and $\mu$. For $\pi^{*}$ a minimizer of (2.1), the 2-Wasserstein metric between $\rho$ and $\mu$ is
$W_{2}(\rho,\mu)=\Bigl{(}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi^{*}(x,y)\Bigr{)}^{1/2}.$ (2.2)
Now, we endow the set of probability measures $\mathcal{P}(\mathcal{X})$ with the 2-Wasserstein distance $W_{2}$. In this paper, we shall represent the set $(\mu^{(i)})_{1\leq i\leq N}$ as the discrete empirical probability measure $\mathbb{P}^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mu^{(i)}}$ over $\mathcal{P}(\mathcal{X})$. To define a metric on $\mathcal{P}(\mathcal{P}(\mathcal{X}))$, the set of Borel probability measures over $\mathcal{P}(\mathcal{X})$, we will use $W_{2}^{2}$ as the ground cost on the metric space $(\mathcal{P}(\mathcal{X}),W_{2})$. The 2-Wasserstein distance over $\mathcal{P}(\mathcal{P}(\mathcal{X}))$ is then defined as [31]
$\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})=\Bigl{(}\min\limits_{\gamma\in\Gamma(\mathbb{P},\mathbb{Q})}\int_{\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})}W_{2}^{2}(\rho,\mu)\mathrm{d}\gamma(\rho,\mu)\Bigr{)}^{1/2},$ (2.3)
where $\Gamma(\mathbb{P},\mathbb{Q})$ is the set of probability distributions on $\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})$ with respective marginals $\mathbb{P}$ and $\mathbb{Q}$.
LOT and KME embeddings.
Given an absolutely continuous (a.c.) measure $\rho$, we first recall that Brenier’s theorem [9] states that the optimal transport plan $\pi^{*}$ in (2.1) is supported on the graph of a $\rho$-a.s. unique push-forward map [^1] We recall that the pushforward of a measure $\rho$ in $\mathbb{R}^{d}$ by a measurable map $T$ is defined as the measure $T_{\#}\rho$ such that for all Borelian $B\subset\mathbb{R}^{d},T_{\#}\rho=\rho\left(T^{-1}(B)\right)$. $T_{\rho}^{\mu}:\mathcal{X}\rightarrow\mathbb{R}^{d}$. In other words, $\pi^{*}=(\mathrm{id},T_{\rho}^{\mu})_{\#}\rho$ and
$W_{2}^{2}(\rho,\mu)=\int_{\mathcal{X}}\|x-T_{\rho}^{\mu}(x)\|^{2}\mathrm{d}\rho(x).$ (2.4)
The LOT embedding then consists in mapping a probability measure $\mu$ to the function $T_{\rho}^{\mu}-\mathrm{id}$ which belongs to the Hilbert space $L^{2}({\rho},\mathbb{R}^{d})=\{v:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}~{}|~{}\int_{\mathbb{R}^{d}}\|v\|^{2}\mathrm{d}{\rho}<\infty\}$, endowed with the weighted $L^{2}$ inner product $\langle v_{1},v_{2}\rangle_{L^{2}({\rho})}=\int_{\mathbb{R}^{d}}v_{1}(x)^{T}v_{2}(x)\mathrm{d}{\rho}(x)$.
Now, given a positive definite kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and associated RKHS $\mathcal{H}$, the KME of $\mu\in\mathcal{P}(\mathcal{X})$ is the embedding $\phi:\mathcal{P}(\mathcal{X})\rightarrow\mathcal{H}$ defined by
$\phi(\mu)=\int_{\mathcal{X}}k(x,\cdot)\mathrm{d}\mathbb{\mu}(x).$ (2.5)
When the kernel $k$ is characteristic [38], the map $\phi$ is injective and one can define a metric on $\mathcal{P}(\mathcal{X})$ called Maximum Mean Discrepancy $\mathrm{MMD}(\rho,\mu)=\|\phi(\rho)-\phi(\mu)\|_{\mathcal{H}}.$
Optimal quantization.
We conclude this section with reminders on the theory of quantization [23, 39]. The optimal quantization of an arbitrary probability measure $\mu$ consists in approximating $\mu$ by a discrete measure that solves the following problem [23][Lemma 3.4]:
$\min\limits_{a\in\Sigma_{K},X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)},$ (2.6)
where $\Sigma_{K}$ is the probability simplex in $\mathbb{R}^{K}$ and $X=(x_{1},\cdots,x_{K})$, with $x_{k}\in\mathbb{R}^{d}$ for all $1\leq k\leq K$. For minimizers $a^{*}$ and $X^{*}$, a $K$-points quantization of $\mu$ is then defined by the discrete measure $\sum_{k=1}^{K}a^{*}_{k}\delta_{x^{*}_{k}}.$
Remark 2.1.
If $\mu$ is a.c., it follows from [32][Proposition 2] that minimizers of (2.6) over $X\in(\mathbb{R}^{d})^{K}$ exist and belong to the set of pairwise distinct points
$F_{K}=\{X=(x_{1},\cdots,x_{K})\in(\mathbb{R}^{d})^{K}~{}|~{}x_{k}\neq x_{\ell},\mbox{ if }k\neq\ell\}.$ (2.7)
Given a vector $X\in F_{K}$, it is well-known, see e.g. [29, 22], that the minimizer $a^{*}$ of $\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ is unique, and verifies $a_{k}^{*}=\mu(V_{x_{k}})$ where $(V_{x_{k}})_{k=1}^{K}$ is the set of Voronoï cells induced by $X$:
$V_{x_{k}}=\bigl{\{}y\in\mathbb{R}^{d}~{}|~{}\forall\ell\neq k,\|x_{k}-y\|^{2}\leq\|x_{\ell}-y\|^{2}\bigr{\}}.$ (2.8)
Thereby, the quantization problem (2.6) rewrites as:
$\min\limits_{X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(V_{x_{k}})\delta_{x_{k}}\Bigr{)}.$ (2.9)
Remark 2.2.
If the measure $\mu$ is discrete, then the closed form solution in variable $a$ given in (2.9) remains valid provided that the definition (2.8) of the Voronoï cells is slightly modified as follows
Ux1:=Vx1and
[…]
so that $(U_{x_{k}})_{k=1}^{K}$ form a partition of $\mathbb{R}^{d}$ [23][Chapter 1] for $X=(x_{1},\ldots,x_{K})\in F_{K}$. This modified definition is needed as some data points might end up on the boundaries of the Voronoï cells (2.8). In that case, the boundaries will have strictly positive measure and the quantization problem (2.6) can no longer be written in the closed-form solution (2.9) from Remark 2.1, that is when $\mu$ is a.c. For such a partitioning (2.10), it follows from Lemma A.1 in Appendix A that
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(U_{x_{k}})\delta_{x_{k}}\Bigr{)}=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y),$
for any probability measure $\mu$. As a consequence, and unless otherwise stated, the results of the paper also hold for discrete probability measures with the choice (2.10) as a Voronoï partition, which corresponds to a chosen enumeration order of the elements of the vector $X$. However, when $\mu$ is a discrete measure, it is necessary to require the cardinality of its support to be larger than $K$ so that the minimizers of (2.6) belong to $F_{K}$ [32][Proposition 2], and the Voronoï cells in the partition (2.10) are pairwise distinct.
Given a $K$-points quantization, that is a minimizer $X^{*}\in(\mathbb{R}^{d})^{K}$ of (2.6), the quantization error of the probability measure $\mu$ is defined as
$\varepsilon_{K}(\mu)=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x^{*}_{k}-y\|^{2}\}\mathrm{d}\mu(y).$ (2.11)
Theorem 6.2 in [23] then implies that $\varepsilon_{K}(\mu)=\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right)$ if $\mu$ is a.c. and $\varepsilon_{K}(\mu)=\mathop{}\mathopen{}o\mathopen{}\left(K^{-2/d}\right)$ if $\mu$ is discrete.
3 Consistency of measure quantizations
Figure 1: Illustration of the quantization methods on two (red and blue) 1D probability measures (left) with $K=2$. (1) Quantization of each measure. (2a) Quantization of the mean measure. (2b) Computation of the weights for each measure. The vertical lines represent the $K=2$ Dirac locations and weights for the quantized measures.
Throughout this section, the supports of the probability measures $(\mu^{(i)})_{i=1}^{N}$ are supposed to be included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$. To reduce the computational costs of solving ML problems involving $N$ measures $(\mu^{(i)})_{i=1}^{N}$ with large supports, we propose two quantization methods to approximate the $\mu^{(i)}$’s with discrete measures supported on $K$ points. These methods are described in the following and illustrated in Figure 1.
3.1 Two quantization schemes
Optimal quantization of each input measure
A first natural approach is to approximate each $\mu^{(i)}$ by its optimal quantization:
$\tilde{\nu}^{(i)}_{K}=\sum\limits_{k=1}^{K}\tilde{a}_{k}^{(i)}\delta_{\tilde{x}_{k}^{(i)}},$ (3.1)
where the weights $\tilde{a}^{(i)}=(\tilde{a}_{k}^{(i)})_{1\leq k\leq K}$ and locations $\widetilde{X}^{(i)}=(\tilde{x}_{k}^{(i)})_{1\leq k\leq K}$ are minimizers of (2.6) for $\mu=\mu^{(i)}$.
Mean-measure quantization
Our second quantization method consists in solving the following problem:
$\min\limits_{a\in(\Sigma_{K})^{N},\ X\in F_{K}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}},\mu^{(i)}\Bigr{)},$ (3.2)
as introduced in [20] for a.c. probability distributions. The following result shows that optimizing (3.2) is equivalent to $K$-points quantization of the mean measure.
Proposition 3.1.
Let $(\mu^{(i)})_{1\leq i\leq N}$ be arbitrary probability measures with support included in a compact set $\mathcal{X}\subset\mathbb{R}^{d}$ and let $\overline{\mu}=\frac{1}{N}\sum_{i=1}^{N}\mu^{(i)}$ be the mean measure. Suppose that the cardinality of the support of $\overline{\mu}$ is larger than $K$. Then,
$\min\limits_{a\in(\Sigma_{K})^{N},\ X\in F_{K}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}},\mu^{(i)}\Bigr{)}$ $=\min\limits_{X\in{(\mathbb{R}^{d})^{K}}}$ $\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}\mu^{(i)}(U_{x_{k}})\delta_{x_{k}},\mu^{(i)}\Bigr{)}$ (3.3)
$=\min\limits_{X\in(\mathbb{R}^{d})^{K}}$ $W_{2}^{2}\Bigl{(}\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}},\overline{\mu}\Bigr{)}$
For a minimizer $\bar{X}=(\bar{x}_{1},\cdots,\bar{x}_{K})$ of (3.3), that is a $K$-point quantization of $\overline{\mu}$, we then define the quantized measures for $1\leq i\leq N$ by
$\bar{\nu}^{(i)}_{K}=\sum\limits_{k=1}^{K}\bar{a}^{(i)}_{k}\delta_{\bar{x}_{k}},\mbox{ with }\bar{a}^{(i)}_{k}=\mu^{(i)}(U_{\bar{x}_{k}}),$ (3.4)
where $(U_{\bar{x}_{k}})_{1\leq k\leq K}$ is the Voronoï partition (2.10) associated to $\bar{X}$. The measure $\bar{\nu}^{(i)}_{K}$ is therefore a discrete probability measure supported on $K$ points that is an approximation of $\mu^{(i)}$ in the sense of the minimization problem (3.2). The measures $(\bar{\nu}^{(i)}_{K})_{1\leq i\leq N}$ differ in their weights but share the same support $\bar{X}$. In a slight abuse of language, we will refer to $\bar{\nu}^{(i)}_{K}$ as a quantized version of $\mu^{(i)}$, even though it is not the optimal quantization $\tilde{\nu}^{(i)}_{K}$ of $\mu^{(i)}$ given in (3.1).
Remark 3.2 (On the compactness assumption of $\mathcal{X}$).
Proposition 3.1 remains true without the compactness assumption on $\mathcal{X}$, under finite $2$-order moments of the measures.
Remark 3.3 (On the nature of input probability measures).
$(i)$ If all the measures $(\mu^{(i)})_{1\leq i\leq N}$ are a.c. then $\overline{\mu}$ is also a.c. and by convention the cardinality of its support is +$\infty$+ ∞. In this case, for all $i\in\{1,\ldots,N\}$ and $k\in\{1,\ldots,K\}$,
$\bar{a}^{(i)}_{k}=\mu^{(i)}(U_{\bar{x}_{k}})=\mu^{(i)}(V_{\bar{x}_{k}}),$
as the boundaries of the Voronoï cells $(V_{\bar{x}_{k}})_{1\leq k\leq K}$ have zero-mass for the Lebesgue measure.
$(ii)$ If all the measures $(\mu^{(i)})_{1\leq i\leq N}$ are discrete, then the definition (3.4) of the probability measure $(\bar{\nu}^{(i)}_{K})_{1\leq i\leq N}$ is specific to the chosen Voronoï partition (2.10) associated to an enumeration of $\bar{X}$. In this setting, a minimizer $\bar{a}$ of (3.2) is not necessarily unique, and another enumeration order of $\bar{X}$ may lead to a slightly different set of quantized measures, depending on the intersection between the points clouds and the boundaries of the Voronoï cells.
For clarity, we write the Voronoï cells associated to a $K$-quantization $\bar{X}$ of the mean measure $\overline{\mu}$ in Prop.3.1 as
$V_{k}:=U_{\bar{x}_{k}},\quad\mbox{for all}\ 1\leq k\leq K.$ (3.5)
3.2 Main result
The following result shows the consistency of both quantization methods by leveraging the quantization error function $\varepsilon_{K}(\cdot)$ defined in (2.11).
Theorem 3.4.
[…]
$\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})=\varepsilon_{K}(\overline{\mu})$ (3.6)
and
$\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)}).$ (3.7)
Remark 3.5.
Theorem 3.4 and the consistency results (3.6) and (3.7) could be extended for more general cost functions than the squared Euclidean cost in the $2$-Wasserstein distance (2.2). For example, for a smooth cost verifying the so-called $x$-regularity from [26][Definition 1], the results would hold.
In other words, Theorem 3.4 shows the convergence of the empirical measures ¯$\overline{\mathbb{P}}^{N}_{K}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}$ towards $\mathbb{P}^{N}$ at the rate $\mathop{}\mathopen{}O\mathopen{}\left(K^{-2/d}\right)$ as $K$ goes to +$\infty$+ ∞.
Remark 3.6.
By definition of optimal quantization (2.6), one has the inequality $W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})$ for all $1\leq i\leq N$. Therefore, we deduce from Theorem 3.4 that $\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})\leq\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$. Hence, ~$\widetilde{\mathbb{P}}^{N}_{K}$ is a better approximation of $\mathbb{P}^{N}$ than ¯$\overline{\mathbb{P}}^{N}_{K}$. Still, the rates of convergence of $\mathcal{W}_{2}^{2}(\widetilde{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ and $\mathcal{W}_{2}^{2}(\overline{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})$ are both scaling as $O(K^{-2/d})$. Moreover, when the $\mu^{(i)}$’s are discrete, mean-measure quantization has computational advantages over optimal quantization of each input measure for moderate to large values of $N$ (see Section 5.1).
Remark 3.7.
In the following, as both quantization methods exhibit similar behavior, we simplify notation by denoting both $\tilde{\nu}^{(i)}_{K}$ (resp. ~$\widetilde{\mathbb{P}}^{N}_{K}$), defined in (3.1), and $\bar{\nu}^{(i)}_{K}$ (resp. ¯$\overline{\mathbb{P}}^{N}_{K}$) defined in (3.4), with a single notation $\nu^{(i)}_{K}$ (resp. $\mathbb{P}^{N}_{K}$). We also write $\varepsilon_{K}=\mathcal{W}_{2}^{2}(\mathbb{P}_{K}^{N},\mathbb{P}^{N})$, where $\varepsilon_{K}=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)})$ in the case of optimal quantization of each input measure (see (3.7)), and $\varepsilon_{K}=\varepsilon_{K}(\overline{\mu})$ in the case of mean-measure quantization (see (3.6)).
Since $\mathcal{X}$ is a compact set, so is the metric space $(\mathcal{P}(\mathcal{X}),W_{2})$ [51][Remark 6.17]. Then, by [45][Theorem 5.9], $\mathcal{W}_{2}(\mathbb{P}^{N}_{K},\mathbb{P})\rightarrow 0$ if and only if $\mathbb{P}^{N}_{K}\rightarrow\mathbb{P}^{N}$ in the sense of weak convergence of distributions, or in other words for any bounded continuous function $f:\mathcal{P}(\mathcal{X})\rightarrow\mathbb{R}$, it holds that $\int f(\nu)\mathrm{d}\mathbb{P}^{N}_{K}(\nu)\overset{K\to+\infty}{\longrightarrow}\int f(\mu)\mathrm{d}\mathbb{P}^{N}(\mu)$. Therefore, one can deduce from Theorem 3.4 the consistency of numerous statistics computed from the quantized measures $(\nu^{(i)}_{K})_{1\leq i\leq N}$ as well as convergence in the MMD sense.
Corollary 3.8.
Let $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ be a positive definite kernel. For any probability measures $\mu,\nu\in\mathcal{P}(\mathcal{X})$, we write $\mathrm{MMD}_{k}(\mu,\nu)=\|\phi(\mu)-\phi(\nu)\|_{\mathcal{H}}$, where $\phi$ and $\mathcal{H}$ are respectively the kernel mean embedding defined in (2.5) and then RKHS associated to $k$. Assume that there exists a constant $C>0$ such that, for all $x,y\in\mathcal{X}$,
$k(x,x)+k(y,y)-2k(x,y)\leq C^{2}\|x-y\|^{2}.$ (3.8)
Then, one has that
$\frac{1}{N}\sum_{i=1}^{N}\mathrm{MMD}_{k}^{2}(\mu^{(i)},\nu^{(i)}_{K})\leq C^{2}\varepsilon_{K}\overset{K\rightarrow\infty}{\longrightarrow}0.$
Corollary 3 in [50] gives mild assumptions for condition (3.8) to hold. It is in particular true for the popular Gaussian kernel.
3.3 Consistent estimation from empirical measures
For an arbitrary measure $\mu$, let us denote $\hat{\mu}=\frac{1}{m}\sum_{j=1}^{m}\delta_{x_{j}}$ the usual empirical measure for $x_{1},\ldots,x_{m}\overset{i.i.d.}{\sim}\mu$. In the above section, we assumed access to the true target measures $(\mu^{(i)})_{i=1}^{N}$. However, in practice when $\mu^{(i)}$ is a.c., we only have access to $\mu^{(i)}$ via random samples, and we can thus only perform quantization from the empirical measure $\hat{\mu}^{(i)}$. Let us denote ^$\hat{\mathbb{P}}_{K}^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\hat{\nu}^{(i)}_{K}}$, where each $\hat{\nu}^{(i)}_{K}$ is obtained by optimal quantization of the input measure $\hat{\mu}^{(i)}$ as described in equation (3.1) in Section 3.1. The following result shows convergence in high probability of ^$\hat{\mathbb{P}}_{K}^{N}$ towards $\mathbb{P}^{N}$.
Theorem 3.9.
Let $(\mu^{(i)})_{1\leq i\leq N}$ be $N$ probability measures supported on a compact set $\mathcal{X}\subset\mathbb{R}^{d}$ with absolutely continuous parts $f^{(i)}\neq 0$. For $\delta>0$, all sufficiently large values of $(m_{i})_{1\leq i\leq N}$ and $K=\max\limits_{1\leq i\leq N}C\cdot m(f^{(i)})\cdot m_{i}^{d/(2d+4)}$, it holds that:
$\mathcal{W}_{2}^{2}(\hat{\mathbb{P}}^{N}_{K},\mathbb{P}^{N})\leq C^{2}\frac{\delta^{2}+\log(N)}{N}\sum\limits_{i=1}^{N}m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}~{}\text{with probability}~{}1-e^{-\delta^{2}},$
where $C$ is a constant that only depends on the dimension $d$ and $m(f)=\int_{\mathcal{X}}f(x)^{d/(d+2)}\mathrm{d}\lambda_{\mathcal{X}}(x)$ with $\lambda_{\mathcal{X}}$ the Lebesgue measure on $\mathcal{X}$.
Note that Theorem 3.9 only holds for quantized measures obtained from the process of optimal quantization of each input measure. Indeed, the arguments used in the proof of Theorem 3.9 cannot be applied when the quantized measures are obtained by mean-measure quantization.
4 Statistics from the quantized measures
We focus here on how statistics computed from the quantized measures $(\nu^{(i)}_{K})_{i=1}^{N}$, defined either by (3.1) or (3.4), relate to statistics computed from the input measures $(\mu^{(i)})_{i=1}^{N}$.
4.1 Wasserstein barycenter
A first example consists in proving that a Wasserstein barycenter [1] of the $(\nu^{(i)}_{K})_{1\leq i\leq N}$ converges towards the unique Wasserstein barycenter of the measures $(\mu^{(i)})_{1\leq i\leq N}$ when at least one of them is a.c.
Proposition 4.1.
Let $\nu_{K}^{\mathrm{bar}}$ be a Wasserstein barycenter of $(\nu^{(i)}_{K})_{1\leq i\leq N}$ that is
$\nu_{K}^{\mathrm{bar}}\in\operatorname*{argmin}\limits_{\nu\in\mathcal{P}(\mathcal{X})}\ \frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\nu,\nu^{(i)}_{K}).$
If at least one of the measures $(\mu^{(i)})_{1\leq i\leq N}$ is a.c., then $\nu_{K}^{\mathrm{bar}}$ converges to the unique Wasserstein barycenter $\mu^{\mathrm{bar}}$ of $(\mu^{(i)})_{1\leq i\leq N}$ in the Wasserstein sense as $K\to+\infty$
4.2 Statistical dispersion
For a set of measures $\mu=(\mu^{(i)})_{i=1}^{N}$, we define its dispersion as the sum of squares $\mathrm{SS}(\mu)=\frac{1}{N^{2}}\sum_{i,j=1}^{N}W_{2}^{2}(\mu^{(i)},\mu^{(j)}).$ The following shows that $\mathrm{SS}(\nu_{K})$, for $\nu_{K}:=(\nu^{(i)}_{K})_{i=1}^{N}$, is controlled by $\mathrm{SS}(\mu)$ and the quantization error $\varepsilon_{K}$ defined in Remark 3.7.
Proposition 4.2.
One has that for any $\lambda>0$,
$\mathrm{SS}(\nu_{K})\leq(1+2/\lambda)\mathrm{SS}(\mu)+(4+2\lambda)\varepsilon_{K}.$
Guaranteeing that the pairwise distance $W_{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})$ is a good approximation of $W_{2}(\mu^{(i)},\mu^{(j)})$ is essential as many machine learning tasks rely on comparing pairs of data. For mean-measure quantization (3.2), we provide below a result on pairwise distances when the input measures are all a.c.
Proposition 4.3.
Suppose that the probability measures $(\mu^{(i)})_{i=1}^{N}$ are a.c. Then, one has for $1\leq i,j\leq N$ and the mean-measure quantization approach in equation (3.4)
$W_{2}^{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})\leq 3W_{2}^{2}(\mu^{(i)},\mu^{(j)})+6\max\limits_{1\leq k\leq K}\mathrm{diam}(V_{k}),$
with $\mathrm{diam}(V_{k})=\max\limits_{x,y\in V_{k}}\|x-y\|^{2}$ and $(V_{k})_{k=1}^{K}$ the Voronoï cells (3.5) obtained from the $K$-points quantization of $\bar{\mu}$.
The following lemma provides an upper bound on the term $\max_{k}\mathrm{diam}(V_{k})$ in Proposition 4.3 in the special case where the support of $\bar{\mu}$ is included in $[0,1]^{d}$. This bound depends on the number of centers $K$ and the ambient dimension $d$, and holds true for either discrete or continuous support.
Lemma 4.4.
Suppose that the (discrete or continuous) support of the mean measure $\bar{\mu}$ is included in $[0,1]^{d}$ and let $(V_{k})_{k=1}^{K}$ be the Voronoï cells of the quantization of $\bar{\mu}$. Then,
$\max\limits_{1\leq k\leq K}\mathrm{diam}(V_{k})\leq\frac{d}{\lfloor\sqrt[d]{K}\rfloor^{2}}.$
4.3 Clustering performances
We now show that both quantization methods preserve the clustering structure of the input measures. To this end, let us assume that each measure $\mu^{(i)}$ has a label $1\leq l\leq L$. We note $I_{l}$ the set of indices such that $\forall i\in I_{l},\mu^{(i)}$ has label $l$, and $N_{l}$ its cardinal. When clustering data, one usually aims at minimizing the within-class variance WCSS for a cluster $l$ and maximizing the between-class variance BCSS for clusters $l_{1}$ and $l_{2}$, where for a set of measure $\mu=(\mu^{(i)})_{i=1}^{N}$,
$\mathrm{WCSS}(l,\mu)$ $=$ $\frac{1}{N_{l}^{2}}\sum\limits_{i,j\in I_{l}}W_{2}^{2}(\mu^{(i)},\mu^{(j)}),$
$\mathrm{BCSS}(l_{1},l_{2},\mu)$ $=$ $\frac{1}{N_{l_{1}}N_{l_{2}}}\sum_{\begin{subarray}{c}i_{1}\in I_{l_{1}}\\ i_{2}\in I_{l_{2}}\end{subarray}}W_{2}^{2}(\mu^{(i_{1})},\mu^{(i_{2})}).$
The next result gives a bound on clustering performances of the quantized measures, and is illustrated in Section 5.
Proposition 4.5.
For a given class $1\leq l\leq L$, one has
$\mathrm{WCSS}(l,\nu_{K})\leq 3\mathrm{WCSS}(l,\mu)+\frac{6N}{N_{l}}\varepsilon_{K}.$ (4.1)
For two distinct classes $l_{1}$ and $l_{2}$, one has that
$\mathrm{BCSS}(l_{1},l_{2},\nu_{K})\geq\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\Bigl{(}\frac{N}{N_{l_{1}}}+\frac{N}{N_{l_{2}}}\Bigr{)}\varepsilon_{K},$
where $\varepsilon_{K}$ is the quantization error defined in Remark 3.7.
4.4 Convergence of the Gram matrices of Hilbert space embeddings after quantization
We consider here the embedding of measures into a Hilbert space using either LOT or KME presented in Section 2. Given the embeddings of both the input measures $(\mu^{(i)})_{i=1}^{N}$ and their quantized approximations $(\nu^{(i)}_{K})_{i=1}^{N}$, we focus on comparing their performances on machine learning methods. To this end, we consider the Gram matrices (4.2) of the pairwise inner-products between the set of embedded measures . Indeed, these matrices play a crucial role in various machine learning tasks [25] such as PCA or Linear Discriminant Analysis (LDA), that rely on the diagonalization of the covariance operator of data in a Hilbert space, which is equivalent to diagonalizing the Gram matrix of inner-products as recalled in Appendix B.
In the following, for a given embedding $\phi:\mathcal{P}(\mathcal{X})\to\mathcal{H}$ of probability measures into a Hilbert space $\mathcal{H}$ equipped with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, we will denote $G^{\phi}_{\mu}$ and $G^{\phi}_{\nu_{K}}$ the $N\times N$ Gram matrices associated with $(\mu^{(i)})_{1\leq i\leq N}$ and $(\nu^{(i)}_{K})_{1\leq i\leq N}$ respectively, with entries
$(G^{\phi}_{\mu})_{ij}=\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}},\mbox{and}\quad(G^{\phi}_{\nu_{K}})_{ij}=\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}.$ (4.2)
We also note $\|\cdot\|_{F}$ the Frobenius matrix norm.
Proposition 4.6.
For $\varepsilon_{K}$ given in Remark 3.7, we have:
(i) We denote $G^{\mathrm{LOT}}_{\mu}$ and $G^{\mathrm{LOT}}_{\nu_{K}}$ the Gram matrices corresponding to the LOT embedding $\phi:\sigma\mapsto T_{\rho}^{\sigma}-\mathrm{id}$, where $\rho\in\mathcal{P}(\mathcal{X})$ is any a.c. reference measure with support included in the compact set $\mathcal{X}\subset\mathbb{R}^{d}$. Assume that the Brenier maps $T_{\rho}^{\mu^{(i)}}$ are $L$-Lipschitz for all $1\leq i\leq N$. Then, we have that
$\frac{1}{N}\|G^{\mathrm{LOT}}_{\mu}-G^{\mathrm{LOT}}_{\nu_{K}}\|_{F}^{2}\leq C_{N,\mathcal{X},L}\sqrt{\varepsilon_{K}}$ (4.3)
where $C_{N,\mathcal{X},L}$ is a constant depending on $N$, the set $\mathcal{X}$ and the Lipschitz constant $L$.
(ii) We note $G^{\mathrm{KME}}_{\mu}$ and $G^{\mathrm{KME}}_{\nu_{K}}$ the Gram matrices corresponding to the KME $\phi:\sigma\mapsto\int k(x,\cdot)\mathrm{d}\sigma(x)$ for a kernel function $k$. Assume that there exists a constant $C>0$ such that, for all $x,y\in\mathcal{X}$,
$k(x,x)+k(y,y)-2k(x,y)\leq C^{2}\|x-y\|^{2}.$
Assume also that $k$ is bounded by a constant $M_{k}<\infty$. Then,
$\frac{1}{N}\|G^{\mathrm{KME}}_{\mu}-G^{\mathrm{KME}}_{\nu_{K}}\|_{F}^{2}\leq C_{N,k}\varepsilon_{K}$ (4.4)
where $C_{N,k}$ is a constant depending on $N$, $M_{k}$ and $C$.
As a consequence of Proposition 4.6, we have that a functional PCA of the maps $(\phi(\nu^{(i)}_{K}))_{i=1}^{N}$ in a certain Hilbert space is consistent (as $K\to+\infty$) with the PCA of the maps $(\phi(\mu^{(i)}))_{i=1}^{N}$ in the same Hilbert space.
4.5 Convergence of PCA in the Wasserstein sense
In this section, we fix a reference measure $\rho$ and we focus on the LOT embedding $\mu\mapsto T_{\rho}^{\mu}-id\in L^{2}(\rho)$. We note $T^{(i)}$ the Monge map between $\rho$ and $\mu^{(i)}$ and $T^{(i)}_{K}$ the Monge map between $\rho$ and $\nu^{(i)}_{K}$ obtained from $\mu^{(i)}$ with any of the two quantization methods presented in Section 3.1. To simplify notation, we note $\overline{T}=T-id$ for a Monge map $T$. In order to define the reconstuction error of PCA, we first define the covariance operators:
$\Sigma=\frac{1}{N}\sum\limits_{i=1}^{N}\overline{T}^{(i)}\otimes\overline{T}^{(i)},\qquad\mbox{and}\quad\Sigma_{K}=\frac{1}{N}\sum\limits_{i=1}^{N}\overline{T}^{(i)}_{K}\otimes\overline{T}^{(i)}_{K},$
where for all $f,g,h\in L^{2}(\rho),(f\otimes g)h=\langle f,h\rangle_{L^{2}(\rho)}g$. Both covariance operators admit spectral representations
$\Sigma=\sum\limits_{j\geq 0}\lambda_{j}P_{j},\qquad\mbox{ and}\quad\Sigma_{K}=\sum\limits_{j\geq 0}(\lambda_{K})_{j}(P_{K})_{j},$
where $(\lambda_{j})_{j\geq 0}$ and $((\lambda_{K})_{j})_{j\geq 0}$ are positive eigenvalues sorted in decreasing order, and the $P_{j}$’s and $(P_{K})_{j}$’s are rank one projectors. Borrowing notation from [43] that is focused in deriving non-asymptotic bounds on the reconstruction error of PCA in a Hilbert space, we write
$P^{\leq q}=\sum\limits_{j\leq q}P_{j},\qquad\mbox{and}\quad P_{K}^{\leq q}=\sum\limits_{j\leq q}(P_{K})_{j}$
the orthogonal projections onto the linear subspaces spanned by the first $q$ eigenvectors of $\Sigma$ and $\Sigma_{K}$. Let $\mathcal{P}_{q}=\{P:L^{2}(\rho)\rightarrow L^{2}(\rho)~{}|~{}P\text{~{}is orthogonal of rank}~{}q\}$. Given $P\in\mathcal{P}_{q}$, we define the reconstruction errors of PCA as:
$Q(P)=\frac{1}{N}\sum\limits_{i=1}^{N}\|\overline{T}^{(i)}-P\overline{T}^{(i)}\|_{L^{2}(\rho)}^{2}\qquad Q_{K}(P)=\frac{1}{N}\sum\limits_{i=1}^{N}\|\overline{T}_{K}^{(i)}-P\overline{T}_{K}^{(i)}\|_{L^{2}(\rho)}^{2}$
As remark in [43] it is well known that PCA amounts to solve
$P^{\leq q}\in\operatorname*{argmin}\limits_{P\in\mathcal{P}_{q}}Q(P)=\operatorname*{argmin}_{P\in\mathcal{P}_{q}}E(P),\quad\mbox{with}\quad E(P)=-\langle\Sigma,P\rangle_{HS},$
where $\langle\cdot,\cdot\rangle_{HS}$ is the Hilbert-Schmidt scalar product defined for linear operators $S,T:L^{2}(\rho)\rightarrow L^{2}(\rho)$ as $\langle S,T\rangle_{HS}=tr(S^{*}T)$. In the same way, we have that:
$P_{K}^{\leq q}\in\operatorname*{argmin}\limits_{P\in\mathcal{P}_{q}}E_{K}(P),\quad\mbox{ where }E_{K}(P)=-\langle\Sigma_{K},P\rangle_{HS}.$
We can finally define the excess risk of the PCA projector $P_{K}^{\leq q}$ with:
$\mathcal{E}^{PCA}_{q}=E(P_{K}^{\leq q})-E(P^{\leq q}).$
The following result allows to control the above excess risk with an upper bound depending on the quantization error $\varepsilon_{K}$.
Proposition 4.7.
Suppose that the $T^{(i)}$’s are $L$-Lipschitz for all $1\leq i\leq N$, then
$\mathcal{E}_{q}^{PCA}\leq 8R\sqrt{qL\ \mathrm{diam}(\mathcal{X})}\varepsilon_{K}^{1/4},$
with $R=\max\limits_{x\in\mathcal{X}}\|x\|$.
5 Numerical experiments
In our experiments, we distinguish the two quantization approaches as follows: $\tilde{K}$-LOT and $\tilde{K}$-KME refer to the LOT embedding and KME of $(\tilde{\nu}^{(i)}_{K})_{i=1}^{K}$ (3.1) obtained from the quantization of each $\mu^{(i)}$, while $\overline{K}$-LOT and $\overline{K}$-KME refer to the LOT embedding and KME of $(\overline{\nu}^{(i)}_{K})_{i=1}^{K}$ (3.4) obtained from the mean-measure quantization. Here, we aim to show that our methods enable fast computation of machine learning tasks while preserving the main information of the measures. We demonstrate their effectiveness by comparing our approaches with several methods:
  1. 1.
    KME with RFF. Random Fourier Features (RFF) [42] allow to efficiently approximate the KME by defining a feature map $\tilde{\varphi}(x)\in\mathbb{R}^{s}$ such that $\tilde{\varphi}(x)^{T}\tilde{\varphi}(y)\approx k(x,y)$. More precisely, this feature map is given by:
    $\tilde{\varphi}(x)=\bigl{(}\sin(\omega_{1}^{T}x),\cdots,\sin(\omega_{s/2}^{T}x),\cos(\omega_{1}^{T}x),\cdots,\cos(\omega_{s/2}^{T}x)\bigr{)},$
    where the $\omega_{i}$’s are independently sampled from a distribution related to the kernel $k$. In this framework, the KME therefore becomes the following mapping between probability measures and $\mathbb{R}^{s}$:
    $\tilde{\phi}(\mu)=\int_{\mathcal{X}}\tilde{\varphi}(x)\mathrm{d}\mu(x).$
    In order to fairly compare results, we choose $s=K$. In our experiments, we use the radial basis kernel (RBF) defined as:
    $k(x,y)=\exp\Bigl{(}-\frac{\|x-y\|^{2}}{2\sigma^{2}}\Bigr{)},$
    where $\sigma$ is a bandwidth parameter. The choice $\sigma$ is made via the median heuristic $\sigma^{2}=\mathrm{median}\{\|X^{(i)}_{j}-X^{(i)}_{l}\|^{2}\}.$
  2. 2.
    $K$-Nys-KME. The method proposed in [13], called Nyström Kernel Mean Embedding, proposes to estimate KME with a Nyström approximation using a small subset of size $K$ (echoing the parameter we deal with in quantization) of the data. More precisely, their approach consists in sampling $K$ points $(x_{k})_{1\leq k\leq K}$ from $\mu$ and find the optimal weights $(\alpha_{k})_{1\leq k\leq K}$ such that the estimator
    $\phi_{K-\mathrm{Nys}}(\mu)=\sum\limits_{k=1}^{K}\alpha_{k}k(x_{k},\cdot)$
    approximates at best the true KME $\phi(\mu)=\int_{\mathcal{X}}k(x,\cdot)\mathrm{d}\mu(x)$.
  3. 3.
    Random subset. Rather than selecting $K$ points with quantization, one could randomly sample $K$ points $(x^{(i)},\cdots,x^{(i)}_{K})$ from $\mu^{(i)}$ and construct the corresponding “non-optimal quantized measure” $\nu^{(i)}_{K}=\sum_{k=1}^{K}\mu^{(i)}(V_{x^{(i)}_{k}})\delta_{x^{(i)}_{k}}$ and then compute the LOT or KME embeddings of these discrete measures supported on a small number of points.
5.1 Computational cost
For a discrete measure, we solve the quantization problem (2.9) using Lloyd’s algorithm [33] and an initialization based on $K$-means+++++ + [4]. The time complexity of the Lloyd’s algorithm being linear in the number of data points [24], it follows that for discrete measures $\mu_{1},\ldots,\mu_{N}$ supported on $m_{1},\ldots,m_{N}$ points respectively, the computational cost for constructing the quantized measures $(\nu^{(i)}_{K})_{i=1}^{K}$ by either optimal quantization of each input measure or mean-measure quantization is $O(Kd\sum_{i=1}^{N}m_{i})$. Nevertheless, as the support of $\overline{\mu}$ is very large in applications, the mean measure quantization can be done on $\frac{1}{N}\sum_{i=1}^{N}m_{i}$ points randomly sampled from $\overline{\mu}$. Indeed, we have observed that subsampling the mean-measure does not deteriorate the performances of our experiments compared to using all points. Then, the computation cost of mean-measure quantization becomes $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})$ and is advantageous over the optimal quantization of each input measure.
In order to compute the LOT embedding, we solve the discrete OT problem (2.1) between $m_{0}$ samples $y_{1},\cdots,y_{m_{0}}$ of $\rho$ and $m_{i}$ samples $x^{(i)}_{1},\cdots,x^{(i)}_{m_{i}}$ of $\mu^{(i)}$ using a standard OT solver [40, 18]. Then, as the optimal map $T_{\rho}^{\mu^{(i)}}$ in (2.4) might not exist, it is classical [16] to compute an approximation through barycentric projection. Let $P^{(i)}\in\mathbb{R}^{m_{0}\times m_{i}}$ be an optimal transport plan, solution of (2.1), between discretized $\rho$ and $\mu^{(i)}$. Then, one defines the barycentric projection map as
$T^{(i)}:y_{j}\mapsto m_{0}\sum\limits_{l=1}^{m_{i}}P^{(i)}_{jl}x^{(i)}_{l}.$
The overall computational complexity of $\overline{K}$-LOT when using $m_{0}$ samples from the reference measure $\rho$ is thus $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})+O(N(K+M)Km_{0}\log(K+m_{0}))$ which is significantly smaller than the one of LOT from the raw input measures that scales as $O(\sum_{i=1}^{N}(m_{i}+m_{0})m_{i}m_{0}\log(m_{i}+m_{0}))$.
Similarly, the computational cost of $K$-KME to construct the Gram matrix $G^{\mathrm{KME}}_{\nu_{K}}$ is $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})+O(N^{2}K^{2})$ that is much cheaper than the cost of computing $G^{\mathrm{KME}}_{\mu}$ from the raw data that scales as $O(\sum_{i,j=1}^{N}m_{i}m_{j})$.
5.2 Synthetic dataset : shifts and scalings of a reference measure
We consider input measures that are shifts and scalings of a given a.c. compactly supported measure $\rho$, that is:
$\mu^{(i)}=(\Sigma_{i}^{1/2}\mathrm{id}+b_{i})_{\#}\rho,$ (5.1)
where $\Sigma_{i}\in\mathbb{R}^{d\times d}$ is a positive semi-definite matrix and $b_{i}\in\mathbb{R}^{d}$. We choose $X\sim\rho$ such that $X=R\frac{Z}{\|Z\|}$, with $R\sim\mathrm{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,I_{d})$. In that case, we have an explicit formulation of the pairwise inner-products induced by LOT embedding and KME, see Proposition C.1 in Appendix C. We can therefore exactly compute the so-called true matrix $G_{\mu}^{\phi}$. In order to compare this matrix to the one computed from our quantizations schemes of empirical measures, we numerically sample the distributions $\mu^{(i)}$’s in the following way : for each $1\leq i\leq N$, we first sample $m_{i}$ points $(x_{j})_{1\leq j\leq m_{i}}$ from the measure $\rho$ by sampling $z_{j}\sim\mathcal{N}(0,I_{d})$ and $r_{j}\sim\mathrm{Unif}([0,1])$ and computing $x_{j}=r_{j}\frac{z_{j}}{\|z_{j}\|}$. This allows to sample points from the unit ball in $\mathbb{R}^{d}$. Samples from $\mu^{(i)}$ are then obtained by the pushforward operation in (5.1).
We first compare in Figure 2 the computational costs of both quantization methods and observe that the mean-measure quantization approach is faster. Additionally, we depict in Figure 4 the evolution of $\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}$ for different values of $K$ and $\phi$ denoting either the LOT or the KME embedding, whose convergence was shown in Proposition 4.6. As expected (see Remark 3.6), the quantization of each measure method yields better approximations than mean-measure quantization. In Figure 3 (resp. Figure 5 in Appendix C), we visualize the projections on the first two components of PCA for $K$-LOT (resp. $K$-KME) of both quantization methods and compare them to the PCA computed from the true Gram matrices $G_{\mu}^{\phi}$ computed in Proposition C.1. We observe that even with a small value $K=32$, the PCA visualizations on quantized embedded measures mimic the ones on raw embedded input measures.
Figure 2: Synthetic dataset on shifts and scalings. Evaluation time for the computation of the two quantization steps for $d=2$ and for different values of $K$.
Figure 3: Synthetic dataset on shifts and scalings. Projection of the data (colored dots) onto the first two components of PCA after $\overline{K}$-LOT (top) and $\tilde{K}$-LOT (bottom) and comparison to the LOT PCA (right) computed from the true Gram matrix (see Prop.C.1).
Figure 4: Synthetic dataset on shifts and scalings. Approximation error of the Gram matrices for $\tilde{K}$-LOT, $\overline{K}$-LOT, $\tilde{K}$-KME, $\overline{K}$-KME in function of $K$.
5.3 Flow cytometry dataset
Table 1: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=64$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 94 30
$\tilde{K}$-LOT 100 81 103
Random subset of size $K$ + LOT 100 77 25
Random subset of size $K$ + KME 100 69 101
$\overline{K}$-KME 100 85 105
$\tilde{K}$-KME 100 69 358
KME with RFF 83 52 5035
$K$-Nys-KME 100 73 36
In this section, we use flow cytometry datasets provided in [48] and publicly available in Mendeley Data to illustrate the suitability of our method through a classification task. We have $N=108$ cytometry measures (or point cloud) which come from two different health care centers : Marburg and Dresden. In Marburg, the data consists of diagnostic samples of peripheral blood (pB), healthy bone marrow (BM), or leukemic bone marrow. The Dresden dataset consists of diagnostic samples of peripheral blood and healthy bone marrow. Two types of labels can be distinguished: data are differentiated either by the healthcare centers from which they were analyzed, or by their types (e.g., peripheral blood, healthy bone marrow, or leukemic bone marrow). Each measure contains from 100,000 to 1,000,000 points in dimension $d=10$, which prevents the use of classical OT, that is, without a quantization step. For $K$-LOT, we sample $m_{0}=1000$ points from the reference measure chosen as the uniform measure on $[0,1]^{d}$ and for $K$-KME, we use the Gaussian kernel with bandwidth parameter $\sigma=1$. The embedded measures live in a high-dimensional Hilbert space (e.g. with $K$-LOT and $K=128$, the ambient space is $\mathbb{R}^{1280}$). In order to keep the most relevant information, we perform a 10-components PCA on the embedded data with respect to the proper Hilbert space $\mathcal{H}$. Then, we train two classifiers (one for each type of label) on 75% of the data and test the results on the remaining data with LDA. Accuracy scores are displayed in Table 1 for $K=64$. Additional experiments for different values of $K$ can be found in Appendix C.3. The quantization methods achieve in particular perfect accuracy scores for predicting the health care center (denoted as LAB) in only a few minutes. We observe that $K$-LOT consistently outperforms $K$-KME. More generally, both approaches outperform the other benchmarked methods. KME with RFF notably performs poorly. These results also underscore the relevance of choosing meaningful centroids with quantization rather than relying on random sampling. Comparing the quantization methods, we observe that the classifier performs better on mean-measure quantization than on the quantization of each measure whereas the latter is roughly ten times slower than the former.
Additionally, we visualize the projections of the data on the first components of PCA in Figures 6 and 7 of the Appendix C. In these figures, it is clear that the representations stabilize when $K\geq 32$. This shows that our $K$-quantization step gives good results even for small values of $K$. We also observe that the mean-measure quantization and the quantization of each measure yield similar results.
5.4 Earth image dataset
Table 2: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=64$.
Method Accuracy Time (s)
$\overline{K}$-LOT 89 20
$\tilde{K}$-LOT 88 305
Random subset of size $K$ + LOT 77 35
Random subset of size $K$ + KME 70 8011
$\overline{K}$-KME 68 7958
$\tilde{K}$-KME 68 9246
$K$-Nys-KME 65 189
We perform similar experiments on a set of images provided by the Airbus company. The dataset consists in $N=1000$ images of size $128\times 128$ captured by a SPOT satellite. The images are divided into two categories : those with the presence of a wind turbine and those without, see Figures 8(a) and 8(b) in Appendix C. Each image is viewed as a discrete probability distribution on the RGB space, that is each pixel is represented by a point in $\mathbb{R}^{3}$. The size $N$ of the dataset as well as the number of pixels ($m_{i}=128^{2}$) prevents from directly computing either LOT or KME. We therefore carry out supervised classification from both embeddings and both quantization methods. We implement $K$-LOT with reference measure the uniform measure on $[0,1]^{3}$ sampled on $m_{0}=1000$ points, and $K$-KME with the Gaussian kernel with bandwidth parameter $\sigma=100$. After the embeddings, we perform PCA and retain only the first 10 components, on which we train an LDA classifier using 75% of the data and test it on the remaining 25%. We obtain the accuracies and execution times displayed in Table 2 for $K=64$, with additional experiments in Appendix C.3 for different values of $K$. Both quantization methods achieve similar accuracy results while mean-measure quantization is approximately ten times faster than the quantization of each measure. We continue to observe that LOT outperforms KME, and that applying quantization to KME yields results comparable to those of other KME-based approaches.
6 Conclusion
In this work, we have proposed to handle machine learning tasks of a set of probability distributions by leveraging two different $K$-quantization approaches that both approximate the input distributions at the asymptotic rate $O(K^{-2/d})$. We proved theoretically that mean-measure quantization and quantization of each measure allow the construction of scalable and consistent embeddings of the probability measures into Hilbert spaces, while numerical experiments highlight the efficiency and accuracy of the former.
One limitation of our quantization approach is its sensitivity to the curse of dimensionality‚ a challenge common to many statistical problems in optimal transport. In future works, one could bypass the dependance on the dimension by relying on some notion of intrinsic dimension, as studied in [53].
7 Acknowledgments
This work benefited from financial support from the French government managed by the National Agency for Research (ANR) under the France 2030 program, with the reference ANR-23-PEIA-0004. The authors also gratefully acknowledge another financial support from the ANR (MaSDOL grant ANR-19-CE23-0017).
Appendix A Proofs of the main results
Lemma A.1.
Let $X=(x_{1},\cdots,x_{K})\in F_{K}$ in (2.7) a vector of distinct points and consider the Voronoï partition
Ux1=Vx1 and
Uxk
[…]
 for k2
[…]
where we recall that
$V_{x_{k}}=\bigl{\{}y\in\mathbb{R}^{d}~{}|~{}\forall\ell\neq k,\|x_{k}-y\|^{2}\leq\|x_{\ell}-y\|^{2}\bigr{\}}.$
Then, for any probability measure $\mu\in\mathcal{P}(\mathcal{X})$, one has that
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ $=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=\sum_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu(y)$
$=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y).$
Proof of Lemma A.1.
Let $X=(x_{1},\cdots,x_{K})\in F_{K}$. From the dual formulation of the Kantorovich problem [40][Equation 5.7], we have that, for any $a\in\Sigma_{K}$,
$W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ $=$ $\sup\limits_{\beta\in\mathbb{R}^{K}}\sum\limits_{k=1}^{K}\beta_{k}a_{k}+\int_{\mathbb{R}^{d}}\beta^{c}(y)\mathrm{d}\mu(y)$ (A.1)
$=$ $\sup\limits_{\beta\in\mathbb{R}^{K}}\sum\limits_{k=1}^{K}a_{k}\beta_{k}+\int_{\mathbb{R}^{d}}\left(\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}-\beta_{k}\}\right)\mathrm{d}\mu(y)$
$\geq$ $\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y),$
where the above inequality is obtain by taking $\beta_{k}=0$ for all $1\leq k\leq K$. Then, since $X=(x_{1},\cdots,x_{K})\in F_{K}$, one has that $(U_{x_{k}})_{1\leq k\leq K}$ is a partition of $\mathbb{R}^{d}$, and we may define
[…]
that is a mapping from $\mathbb{R}^{d}$ to $X$. Introducing the probability measure $\mu_{K}=\sum\limits_{k=1}^{K}\mu(U_{x_{k}})\delta_{x_{k}},$ it is not difficult to see that $T_{K\#}\mu=\mu_{K}$ where the notation $T_{\#}\mu$ denotes the push-forward of a measure $\mu$ by the mapping $T$. Now, we let $\pi_{K}=(\mathrm{id}\times T_{K})_{\#}\mu$ that obviously belongs to the set of transport plans $\Pi(\mu,\mu_{K})$. From the definition of the Voronoï partition $(U_{x_{k}})_{1\leq k\leq K}$, one can than check that, for any $y\in\mathbb{R}^{d}$, (see e.g. [39])
$\|y-T_{K}(y)\|^{2}=\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}.$
Consequently, we obtain the following equalities
$\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi_{K}(x,y)=\int_{\mathbb{R}^{d}}\|y-T_{K}(y)\|^{2}\mathrm{d}\mu(y)=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y).$
Inserting the above equality into (A.1), we thus have that, for any $a\in\Sigma_{K}$,
$W_{2}^{2}(\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}})\geq\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi_{K}(x,y).$
Since $W_{2}^{2}\Bigl{(}\mu,\mu_{K}\Bigr{)}=\min\limits_{\pi\in\Pi(\mu,\mu_{K})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi(x,y)$, we directly obtain from the above inequality that $W_{2}^{2}(\mu,\mu_{K})=\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{2}\mathrm{d}\pi_{K}(x,y)$, which implies
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\mu_{K}\Bigr{)}=\int_{\mathbb{R}^{d}}\|y-T_{K}(y)\|^{2}\mathrm{d}\mu(y)=\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\mu(y),$
which concludes the proof. ∎
Proof of Proposition 3.1.
Let $X\in F_{K}$. Applying Lemma A.1, we obtain that, for any $1\leq i\leq N$,
$\min\limits_{a^{(i)}\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu^{(i)},\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu^{(i)}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}=\sum_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu^{(i)}(y),$
and thus we have that
$\min\limits_{a\in(\Sigma_{K})^{N}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\mu^{(i)},\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}}\Bigr{)}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu^{(i)}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu^{(i)}(y)$
$=$ $\sum\limits_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\overline{\mu}(y)$
$=$ $W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=$ $\int_{\mathbb{R}^{d}}\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}\}\mathrm{d}\overline{\mu}(y),$
where we again apply Lemma A.1 to derive the last equality above. Finally, from the assumption that the cardinality of the support of $\overline{\mu}$ is larger than $K$, we obtain from [32][Proposition 2] that
$\min\limits_{X\in F_{K}}W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}=\min\limits_{X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
which concludes the proof.
Proof of Theorem 3.4.
For a fixed $N\geq 1$, since $\mathbb{P}^{N}$, ¯$\overline{\mathbb{P}}^{N}_{K}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}$ are discrete uniform measures of the same size, one can actually restrict the optimization set in (2.3) to permutations $\sigma\in\mathrm{Perm}(N)$ [40][Equation 2.2] in the following sense:
$\mathcal{W}_{2}^{2}(\mathbb{P}^{N},\overline{\mathbb{P}}^{N}_{K})$ $=\min\limits_{\sigma\in\mathrm{Perm}(N)}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(\sigma(i))}_{K})$
$\mathcal{W}_{2}^{2}(\mathbb{P}^{N},\widetilde{\mathbb{P}}^{N}_{K})$ $=\min\limits_{\sigma\in\mathrm{Perm}(N)}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(\sigma(i))}_{K})$ (A.2)
However, for $\tilde{\nu}^{(i)}_{K}$, defined in (3.1), it follows by the definition of optimal quantization of $\mu^{(i)}$ that, for any $1\leq j\leq N$,
$W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(j)}_{K}).$ (A.3)
Similarly, $\bar{\nu}^{(i)}_{K}$ defined in (3.4) corresponds to the discrete probability measure supported on $\bar{X}$ that best approximates $\mu^{(i)}$. In other words, $W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\sum_{k=1}^{K}a_{k}\delta_{\bar{x}_{k}})$ for any weight vector $a\in\Sigma_{K}$. In particular, we have that, for any $1\leq j\leq N$,
$W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(j)}_{K}).$ (A.4)
Using Inequalities (A.3) and (A.4), it is then easy to see that in both cases, the optimal permutation minimizing (A.2) is the identity $\sigma(i)=i$ for all $1\leq i\leq N$, and that we have
$\mathcal{W}_{2}^{2}(\mathbb{P}^{N},\overline{\mathbb{P}}^{N}_{K})=\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\overline{\nu}^{(i)}_{K})\qquad\mathcal{W}_{2}^{2}(\mathbb{P}^{N},\widetilde{\mathbb{P}}^{N}_{K})=\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})$
For $\tilde{\nu}^{(i)}_{K}$, one immediately has that $\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})=\frac{1}{N}\sum_{i=1}^{N}\varepsilon_{K}(\mu^{(i)})$. For $\bar{\nu}^{(i)}_{K}$, we obtain from Proposition 3.1 that
$\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})=W_{2}^{2}\left(\sum\limits_{k=1}^{K}\overline{\mu}(U_{\bar{x}_{k}})\delta_{\bar{x}_{k}},\overline{\mu}\right)=\varepsilon_{K}(\overline{\mu}),$
which concludes the proof. ∎
Proof of Corollary 3.8.
We have that $(\mathcal{X},\|\cdot\|)$ is a complete separable metric space, $k$ is positive definite and verifies $\forall x,y\in\mathcal{X},k(x,x)+k(y,y)-2k(x,y)\leq C^{2}\|x-y\|^{2}$ for a constant $C>0$, and $\mu^{(i)}$ and $\nu^{(i)}$ have finite 2-moments as they are supported on a compact set $\mathcal{X}\subset\mathbb{R}^{d}$. We can then use Proposition 2 of [50], which proves that under these assumptions,
$\mathrm{MMD}(\mu^{(i)},\nu_{K}^{(i)})\leq CW_{2}(\mu^{(i)},\nu_{K}^{(i)}).$
We can directly conclude with Theorem 3.4 that:
$\frac{1}{N}\sum_{i=1}^{N}\text{MMD}^{2}(\mu^{(i)},\nu^{(i)}_{K})$$\leq$ $\frac{1}{N}\sum_{i=1}^{N}C^{2}W_{2}^{2}(\mu^{(i)},\nu^{(i)}_{K})$
$=$ $C^{2}\varepsilon_{K}\overset{K\rightarrow\infty}{\longrightarrow}0.$
Proof of Theorem 3.9.
The proof of Theorem 3.9 relies on [10][Theorem 5.2], that we recall below for completeness.
Theorem A.2 (Theorem 5.2 from [10]).
Let $\mu\in\mathcal{P}(\mathcal{X})$ with absolutely continuous part $f\neq 0$. For $\tau>0$, all sufficiently large values of $m$ and $K=C\cdot M(f)\cdot m^{d/(2d+4)}$, it holds that
$W_{2}(\mu,\hat{\nu}_{K})\leq C\cdot m(f)\cdot m^{-1/(2d+4)\cdot\tau}~{}\text{with probability}~{}1-e^{-\tau^{2}},$
where $\hat{\nu}_{K}$ is the $K$-points optimal quantization of the empirical measure $\hat{\mu}=\frac{1}{m}\sum_{j=1}^{m}\delta_{x_{j}}$ associated to $\mu$ supported on $m$ points and $C$ is a constant that only depends on $d$.
Let $\hat{\nu}^{(i)}_{K^{(i)}}$ be the $K^{(i)}$-points optimal quantization of $\hat{\mu}^{(i)}$, the empirical version of $\mu^{(i)}$ supported on $m_{i}$ points. Adapting Theorem A.2 in this context gives that for $\tau>0$, for sufficiently large $m_{i}$ and letting $K^{(i)}=C\cdot m(f^{(i)})\cdot m_{i}^{d/(2d+4)}$, it holds that
$W^{2}_{2}(\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}})\leq C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}~{}\text{with~{} probability}~{}1-e^{-\tau^{2}},$ (A.5)
In other words, Equation (A.5) reads
$\mathbf{Pr}\Bigl{[}W_{2}^{2}\bigl{(}\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}\leq C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-e^{-\tau^{2}}$
If we denote $A_{i}$ the event $\{W_{2}^{2}\bigl{(}\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}>C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\}$, we obtain that
$\mathbf{Pr}[A_{i}]\leq e^{-\tau^{2}}\Longrightarrow\mathbf{Pr}\Bigl{[}\cup_{i=1}^{N}A_{i}\Bigr{]}\leq\sum\limits_{i=1}^{N}e^{-\tau^{2}}=Ne^{-\tau^{2}},$
which can be rewritten as
$\mathbf{Pr}$ $\Bigl{[}\exists i,W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}>C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\leq Ne^{-\tau^{2}}$
$\Longrightarrow$ $\mathbf{Pr}$ $\Bigl{[}\forall i,W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}\leq C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-Ne^{-\tau^{2}}$
$\Longrightarrow$ $\mathbf{Pr}$ $\Bigl{[}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}\leq\frac{1}{N}\sum\limits_{i=1}^{N}C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-Ne^{-\tau^{2}}$
Now, taking $K=\max\limits_{1\leq i\leq N}K^{(i)}$, one has $\forall 1\leq i\leq N$, $W^{2}_{2}(\mu^{(i)},\hat{\nu}^{(i)}_{K})\leq W^{2}_{2}(\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}})$. Hence, we finally obtain that
$\mathbf{Pr}\Bigl{[}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K}\bigr{)}\leq\frac{1}{N}\sum\limits_{i=1}^{N}C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-Ne^{-\tau^{2}}$
Now taking $\tau=\sqrt{\delta^{2}+\log(N)}$, one has that
$\mathbf{Pr}\Bigl{[}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K}\bigr{)}\leq C^{2}\frac{\delta^{2}+\log(N)}{N}\sum\limits_{i=1}^{N}m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\Bigr{]}\geq 1-e^{-\delta^{2}}.$
We can conclude the proof of Theorem 3.9 by observing that $\mathcal{W}_{2}^{2}(\hat{\mathbb{P}}_{K}^{N},\mathbb{P}^{N})\leq\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\hat{\nu}_{K}^{(i)})$.
Proof of Proposition 4.1.
For a fixed $N$, and thanks to Theorem 3.4, we have that the sequence of probability measures $(\mathbb{P}^{N}_{K})_{K\geq 1}$ such that $\mathbb{P}^{N}_{K}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\nu^{(i)}_{K}}\subset\mathcal{P}(\mathcal{P}(\mathcal{X}))$ converges towards $\mathbb{P}^{N}$, that is $\mathcal{W}_{2}(\mathbb{P}^{N}_{K},\mathbb{P}_{N})\overset{K\rightarrow\infty}{\longrightarrow}0$. Additionally, the Wasserstein barycenter of $\mathbb{P}^{N}$ is unique (Proposition 3.5 in [1]) since at least one of the probability measures $\mu^{(i)},1\leq i\leq N$ is a.c. by hypothesis. Therefore, using [31][Theorem 3], we immediately obtain that the sequence of barycenters of $(\mathbb{P}^{N}_{K})_{K\geq 1}$ converges towards the barycenter of $\mathbb{P}^{N}$ in the $W_{2}$ distance.
Proof of Proposition 4.2.
From the triangle inequality, one can write:
$W_{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq W_{2}(\nu^{(i)}_{K},\mu^{(i)})+W_{2}(\mu^{(i)},\mu^{(j)})+W_{2}(\mu^{(j)},\nu^{(j)}_{K}).$
From Young’s inequality, $2ab\leq a^{2}+b^{2}$ and $2ab\leq\lambda a^{2}+\frac{b^{2}}{\lambda}$ for any $\lambda>0$ and any real numbers $a,b$ and $c$. Then it holds that :
$(a+b+c)^{2}\leq(2+\lambda)a^{2}+(2+\lambda)c^{2}+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}b^{2}$ (A.6)
Squaring the triangle inequality and using (A.6) yields to:
$W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq(2+\lambda)W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}W_{2}^{2}(\mu^{(i)},\mu^{(j)})+(2+\lambda)W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})$ (A.7)
We now sum inequality (A.7) over all $1\leq i,j\leq N$, and divide by $N^{2}$:
$\mathrm{SS}(\nu_{K})$ $=\frac{1}{N^{2}}\sum\limits_{i,j=1}^{N}W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq\frac{2}{N}(2+\lambda)\sum\limits_{i=1}^{N}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\frac{1}{N^{2}}\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}\sum\limits_{i,j=1}^{N}W_{2}^{2}(\mu^{(i)},\mu^{(j)})$
Hence, by Theorem 3.4, we obtain that $\mathrm{SS}(\nu_{K})\leq(4+2\lambda)\varepsilon_{K}+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}\mathrm{SS}(\mu)$, which concludes the proof. ∎
Proof of Proposition 4.3.
We first recall that the dual formulation of OT between the discrete measures $\bar{\nu}^{(i)}_{K}$ and $\bar{\nu}^{(j)}_{K}$ (see e.g. [40]) is given by
$W_{2}^{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})=\max_{(\alpha,\beta)\in\Phi}\ \sum_{k=1}^{K}a_{k}^{(i)}\alpha_{k}+\sum_{k=1}^{K}a_{k}^{(j)}\beta_{k}$ (A.8)
where $\Phi:=\{(\alpha,\beta)\in\mathbb{R}^{K}\times\mathbb{R}^{K}\ \mbox{such that for all}\ 1\leq k,l\leq K,\alpha_{k}+\beta_{l}\leq\|\bar{x}_{k}-\bar{x}_{l}\|^{2}\}$. Let $\alpha^{ij},\beta^{ij}\in\mathbb{R}^{K}$ be optimal Kantorovich potentials for $\bar{\nu}^{(i)}_{K}$ and $\bar{\nu}^{(j)}_{K}$ in (A.8). We define the piecewise constant function $f^{ij}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $x\mapsto\alpha^{ij}_{k}$ when $x\in\textrm{int}(V_{k})$, where $\textrm{int}(V_{k})$ denotes the open interior of the Voronoï cell $V_{k}$. Similarly, $g^{ij}:y\mapsto\beta^{ij}_{k}$ when $y\in V_{k}$. Then, thanks to the absolute continuity of the $\mu^{(i)}$’s, one can write:
$W_{2}^{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})$ $=\sum\limits_{k=1}^{K}a^{(i)}_{k}\alpha^{ij}_{k}+\sum\limits_{k=1}^{K}a^{(j)}_{k}\beta^{ij}_{k}$
$=\sum\limits_{k=1}^{K}\int_{V_{k}}\mathrm{d}\mu^{(i)}(x)\alpha^{ij}_{k}+\sum\limits_{k=1}^{K}\int_{V_{k}}\mathrm{d}\mu^{(j)}(y)\beta^{ij}_{k}$
$=\sum\limits_{k=1}^{K}\int_{\textrm{int}(V_{k})}\alpha^{ij}_{k}\mathrm{d}\mu^{(i)}(x)+\sum\limits_{k=1}^{K}\int_{\textrm{int}(V_{k})}\beta^{ij}_{k}\mathrm{d}\mu^{(j)}(y)$
$=\int_{\mathbb{R}^{d}}f^{ij}(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g^{ij}(y)\mathrm{d}\mu^{(j)}(y)$ (A.9)
We then aim at identifying (A.9) with a dual formulation for OT between $\mu^{(i)}$ and $\mu^{(j)}$ with respect to some cost function $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ to be defined later on, where
$\text{OT}_{c}(\mu^{(i)},\mu^{(j)})$ $=\min\limits_{\pi\in\Pi(\mu^{(i)},\mu^{(j)})}\int c(x,y)\mathrm{d}\pi(x,y)$
$=\sup\limits_{f,g:f(x)+g(y)\leq c(x,y)}\int_{\mathbb{R}^{d}}f(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g(y)\mathrm{d}\mu^{(j)}(y).$ (A.10)
If $x\in V_{k}$ and $y\in V_{k^{\prime}}$, we obtain
$f^{ij}(x)+g^{ij}(y)$ $=\alpha^{ij}_{k}+\beta^{ij}_{k^{\prime}}$
$\leq\|x_{k}-x_{k^{\prime}}\|^{2}$
$\leq 3\|x_{k}-x\|^{2}+3\|x-y\|^{2}+3\|y-x_{k^{\prime}}\|^{2}$
$\leq 3\mathrm{diam}(V_{k})+3\mathrm{diam}(V_{k^{\prime}})+3\|x-y\|^{2}$
$\leq 6\max\limits_{k}\mathrm{diam}(V_{k})+3\|x-y\|^{2},$ (A.11)
where the first inequality is due to the fact that $\alpha^{ij}_{k}$ and $\beta^{ij}_{k^{\prime}}$ are Kantorovich potentials of $W_{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})$ in (A.8). The second inequality comes from (A.6) where $\lambda=1$. Defining the new cost function $c(x,y)=6\max\limits_{k}\mathrm{diam}(V_{k})+3\|x-y\|^{2}$, we thus define
$\text{OT}_{c}(\mu^{(i)},\mu^{(j)})$ $=6\max\limits_{k}\mathrm{diam}(V_{k})+3\min\limits_{\pi\in\Pi(\mu^{(i)},\mu^{(j)})}\int\|x-y\|^{2}\mathrm{d}\pi(x,y)$
$=6\max\limits_{k}\mathrm{diam}(V_{k})+3W_{2}^{2}(\mu^{(i)},\mu^{(j)}).$
Now, by inequality (A.11), it follows that $f^{ij}$ and $g^{ij}$ are feasible Kantorovich potentials of $\text{OT}_{c}$ between $\mu^{(i)}$ and $\mu^{(j)}$ in (A.10). Hence, we finally obtain from (A.9) that
$W_{2}^{2}(\bar{\nu}^{(i)}_{K},\bar{\nu}^{(j)}_{K})$ $\leq\int_{\mathbb{R}^{d}}f^{ij}(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g^{ij}(y)\mathrm{d}\mu^{(j)}(y)$
$\leq\sup\limits_{f,g:f(x)+g(y)\leq c(x,y)}\int_{\mathbb{R}^{d}}f(x)\mathrm{d}\mu^{(i)}(x)+\int_{\mathbb{R}^{d}}g(y)\mathrm{d}\mu^{(j)}(y)$
$=6\max\limits_{k}\mathrm{diam}(V_{k})+3W_{2}^{2}(\mu^{(i)},\mu^{(j)}),$
which concludes the proof. ∎
Proof of Lemma 4.4.
Suppose that the support of the mean measure $\bar{\mu}$ is included in $[0,1]^{d}$. Then, we first have that $\max\limits_{1\leq k\leq K}\mathrm{diam}(V_{k})\leq\max\limits_{1\leq j\leq\lfloor\sqrt[d]{K}\rfloor^{d}}\mathrm{diam}(V_{j})$. Indeed, as $d]{K}\rfloor^{d}\leq K$, this is simply reducing the number of quantization points, and therefore increasing the maximum diameter of the cells. Now, denoting $K^{\prime}=\lfloor\sqrt[d]{K}\rfloor^{d}$ the $d$-th power of the integer $d]{K}\rfloor$, one can grid the support space $[0,1]^{d}$ with $K^{\prime}$ points $\{x_{1},\ldots,x_{K^{\prime}}\}$ set as $\Bigl{\{}\left(\frac{a^{(i)}_{1}}{\lfloor\sqrt[d]{K}\rfloor},\cdots,\frac{a^{(i)}_{d}}{\lfloor\sqrt[d]{K}\rfloor}\right)~{}|~{}a^{(i)}_{k}\in\{1,\cdots,d\}\Bigr{\}}$. With these centers, all Voronoï cells have the same diameter, which is:
[…]
This finally gives us:
$\max\limits_{1\leq k\leq K}\mathrm{diam}(V_{k})\leq\frac{d}{\lfloor\sqrt[d]{K}\rfloor^{2}}.$
Proof of Proposition 4.5.
For the result regarding the within-class variance of the clusters, we have, using the triangle inequality and (A.6) with $\lambda=1$,
$W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq 3\bigl{(}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+W_{2}^{2}(\mu^{(i)},\mu^{(j)})+W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})\bigr{)}$
Summing over the indices of $I_{l}$ and dividing by $N_{l}^{2}$ yields:
$\mathrm{WCSS}(l,\nu_{K})$ $\leq\frac{3}{N_{l}^{2}}\sum\limits_{i,j\in I_{l}}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\frac{3}{N_{l}^{2}}\sum\limits_{i,j\in I_{l}}W_{2}^{2}(\mu^{(i)},\mu^{(j)})+\frac{3}{N_{l}^{2}}\sum\limits_{i,j\in I_{l}}W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})$
$\leq\frac{6}{N_{l}}\sum\limits_{i\in I_{l}}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+3\mathrm{WCSS}(l,\mu)$
$\leq\frac{6}{N_{l}}\sum\limits_{1\leq i\leq N}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+3\mathrm{WCSS}(l,\mu)=\frac{6N}{N_{l}}\varepsilon_{K}+3\mathrm{WCSS}(l,\mu),$
where the last equality follows from Theorem 3.4, which concludes the first item (4.1) of the proposition. For the second statement on the between-class variance, we rewrite the triangle inequality:
$\quad 3\bigl{(}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})+W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})\bigr{)}\geq W_{2}^{2}(\mu^{(i)},\mu^{(j)})$
$\Leftrightarrow W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\geq\frac{1}{3}W_{2}^{2}(\mu^{(i)},\mu^{(j)})-W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})-W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})$
Summing over the indices of $I_{l_{1}}$ and $I_{l_{2}}$ and dividing by $N_{l_{1}}N_{l_{2}}$ gives:
$\mathrm{BCSS}(l_{1},l_{2},\nu_{K})$ $\geq\frac{1}{3}\frac{1}{N_{l_{1}}N_{l_{2}}}\sum\limits_{\begin{subarray}{c}i_{1}\in I_{l_{1}}\\ i_{2}\in I_{l_{2}}\end{subarray}}W_{2}^{2}(\mu^{(i_{1})},\mu^{(i_{2})})$
$\qquad-\frac{1}{N_{l_{1}}N_{l_{2}}}\sum\limits_{\begin{subarray}{c}i_{1}\in I_{l_{1}}\\ i_{2}\in I_{l_{2}}\end{subarray}}W_{2}^{2}(\nu^{(i_{1})}_{K},\mu^{(i_{1})})-\frac{1}{N_{l_{1}}N_{l_{2}}}\sum\limits_{\begin{subarray}{c}i_{1}\in I_{l_{1}}\\ i_{2}\in I_{l_{2}}\end{subarray}}W_{2}^{2}(\mu^{(i_{2})},\nu^{(i_{2})}_{K})$
$\geq\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\frac{1}{N_{l_{1}}N_{l_{2}}}\sum\limits_{1\leq i_{1},i_{2}\leq N}W_{2}^{2}(\nu^{(i_{1})}_{K},\mu^{(i_{1})})$
$\qquad-\frac{1}{N_{l_{1}}N_{l_{2}}}\sum\limits_{1\leq i_{1},i_{2}\leq N}W_{2}^{2}(\nu^{(i_{2})}_{K},\mu^{(i_{2})})$
$=\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\frac{1}{N_{l_{1}}}\sum\limits_{1\leq i_{1}\leq N}W_{2}^{2}(\nu^{(i_{1})}_{K},\mu^{(i_{1})})$
$\qquad-\frac{1}{N_{l_{2}}}\sum\limits_{1\leq i_{2}\leq N}W_{2}^{2}(\nu^{(i_{2})}_{K},\mu^{(i_{2})})$
$=\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\Bigl{(}\frac{1}{N_{l_{1}}}+\frac{1}{N_{l_{2}}}\Bigr{)}\sum\limits_{1\leq i\leq N}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})$
$=\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\Bigl{(}\frac{N}{N_{l_{1}}}+\frac{N}{N_{l_{2}}}\Bigr{)}\varepsilon_{K},$
where the last equality follows from Theorem 3.4, which concludes the proof. ∎
The proof of Proposition 4.6 relies on the following Lemma which holds for any embedding $\phi$.
Lemma A.3.
Let $\phi:\mathcal{P}(\mathcal{X})\rightarrow\mathcal{H}$ be an embedding of probability measures (supported on an arbitrary set included in $\mathcal{X}\subset\mathbb{R}^{d}$) into a Hilbert space $\mathcal{H}$ equipped with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ and the induced norm $\|\cdot\|_{\mathcal{H}}$. Suppose that $\|\phi\|_{\infty}=:M_{\phi}<\infty$. Then, we have that:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}\leq C\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}$ (A.12)
where $C$ is a constant depending on $N$ and $M_{\phi}$.
Proof of Lemma A.3.
First, let us write the Frobenius matrix norm of $G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}$:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}=\sum\limits_{i,j=1}^{N}|(G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}})_{ij}|^{2}=\sum\limits_{i,j=1}^{N}|\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}|^{2}$ (A.13)
Additionally, we have that:
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}-\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$=\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(j}))-\phi(\nu^{(j)}_{K}),\phi(\nu^{(i)}_{K})\rangle_{\mathcal{H}}.$
Injecting this equality in (A.13) yields:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}$ $=\sum\limits_{i,j=1}^{N}\Bigl{|}\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}+\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}$
$\quad+\langle\phi(\mu^{(j}))-\phi(\nu^{(j)}_{K}),\phi(\nu^{(i)}_{K})\rangle_{\mathcal{H}}\Bigr{|}^{2}$
$\leq 3\sum\limits_{i,j=1}^{N}\Bigl{|}\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}\Bigr{|}^{2}$
$\quad+3\sum\limits_{i,j=1}^{N}\Bigl{|}\langle\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}\Bigr{|}^{2}$
$\quad+3\sum\limits_{i,j=1}^{N}\Bigl{|}\langle\phi(\mu^{(j}))-\phi(\nu^{(j)}_{K}),\phi(\nu^{(i)}_{K})\rangle_{\mathcal{H}}\Bigr{|}^{2}$
$\leq 3\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigl{\|}\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}$ (A.14)
$\qquad+3\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\Bigr{\|}^{2}_{\mathcal{H}}\Bigl{\|}\phi(\nu^{(j)}_{K})\Bigr{\|}^{2}_{\mathcal{H}}$
$\qquad+3\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(j)})-\phi(\nu^{(j)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigl{\|}\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}$
$=3\Bigl{(}\sum\limits_{i=1}^{N}\Bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\Bigr{\|}_{\mathcal{H}}^{2}\Bigr{)}^{2}+6\sum\limits_{i,j=1}^{N}\Bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\Bigr{\|}^{2}_{\mathcal{H}}\Bigl{\|}\phi(\nu^{(j)}_{K})\Bigr{\|}^{2}_{\mathcal{H}}$
$\leq\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigl{(}3\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}+6\sum\limits_{j=1}^{N}\bigl{\|}\phi(\nu^{(j)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigr{)}$ (A.15)
where (A.14) is due to Cauchy-Schwarz. By hypothesis, we have $\|\phi\|_{\infty}=\sup\limits_{\mu\in\mathcal{P}(\mathcal{X})}\|\phi(\mu)\|_{\mathcal{H}}=M_{\phi}<\infty$. Therefore, (A.15) yields:
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}$ $\leq\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}\Bigl{(}12\sum\limits_{i=1}^{N}M_{\phi}^{2}+6\sum\limits_{j=1}^{N}M_{\phi}^{2}\Bigr{)}$
$=18NM_{\phi}^{2}\sum\limits_{i=1}^{N}\bigl{\|}\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\bigr{\|}^{2}_{\mathcal{H}}.$
Choosing $C:=18NM_{\phi}^{2}$ concludes the proof.
Proof of Proposition 4.6.
(ii) We first prove inequality (4.4) for the KME embedding $\phi:\sigma\in\mathcal{P}(\mathcal{X})\mapsto\int k(x,\cdot)\mathrm{d}\sigma(x)\in\mathcal{H}$. By hypothesis, the kernel $k$ is bounded by a constant $M_{k}$. Therefore for any $\mu\in\mathcal{P}(\mathcal{X})$, we have $\|\phi(\mu)\|_{\mathcal{H}}^{2}=\langle\int k(x,\cdot)\mathrm{d}\mu(x),\int k(y,\cdot)\mathrm{d}\mu(y)\rangle_{\mathcal{H}}=\iint k(x,y)\mathrm{d}\mu(x)\mathrm{d}\mu(y)\leq M_{k}$. By Lemma A.3, we then have
$\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}^{2}\leq C_{N,k}\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}$
with $C_{N,k}:=18NM_{k}$. As in the proof of Corollary 3.8, under our assumptions, we have that $\mathrm{MMD}_{k}\leq CW_{2}$. we obtain that
$\frac{1}{N}\|G_{\mu}^{\mathrm{KME}}-G_{\nu_{K}}^{\mathrm{KME}}\|_{F}^{2}$ $\leq\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}\|\phi(\mu^{(i)})-\phi(\nu^{(i)}_{K})\|^{2}_{\mathcal{H}}=\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}\mathrm{MMD}^{2}_{k}(\mu^{(i)},\nu^{(i)}_{K})$
$\leq\frac{C_{N,k}}{N}\sum\limits_{i=1}^{N}C^{2}W_{2}^{2}(\mu^{(i)},\nu^{(i)}_{K})$
$=C_{N,k}\ \varepsilon_{K},$
if we redefine $C_{N,k}=18NM_{k}C^{2}$.
(i) In order to prove inequality (4.3) for the LOT embedding $\phi:\mu\in\mathcal{P}(\mathcal{X})\mapsto T_{\rho}^{\mu}-\mathrm{id}\in L^{2}(\rho)$ with a.c. reference measure $\rho$, we first note that for any measure $\mu\in\mathcal{P}(\mathcal{X})$, $\|\phi(\mu)\|^{2}_{L^{2}(\rho)}=\int|T_{\rho}^{\mu}(x)-x|^{2}\mathrm{d}\rho(x)\leq\int\mathrm{diam}(\mathcal{X})^{2}\mathrm{d}\rho(x)=\mathrm{diam}(\mathcal{X})^{2}$. Then we can apply Lemma A.3 with constant $C:=C_{N,\mathcal{X}}=18N\mathrm{diam}(\mathcal{X})^{2}$ in (A.12) and finally obtain
$\frac{1}{N}\|G_{\mu}^{\mathrm{LOT}}-G_{\nu_{K}}^{\mathrm{LOT}}\|_{F}^{2}$ $\leq\frac{C_{N,\mathcal{X}}}{N}\sum\limits_{i=1}^{N}\|T_{\rho}^{\mu^{(i)}}-T_{\rho}^{\nu^{(i)}_{K}}\|_{L^{2}(\rho)}^{2}.$ (A.16)
We now use a result due to Ambrosio and reported in [21] and [17][Theorem (Ambrosio)], which states that when $\rho$ is a probability density over a compact set $\mathcal{X}$, $\mu$ and $\nu$ are probability measures on a $\mathcal{X}$ and $T_{\rho}^{\mu}$ is $L$-Lipschitz (by hypothesis), then:
$\|T_{\rho}^{\mu}-T_{\rho}^{\nu}\|_{L^{2}(\rho)}\leq 2\sqrt{\mathrm{diam}(\mathcal{X})L}W_{1}(\mu,\nu)^{1/2}.$ (A.17)
Finally, assembling the inequalities (A.16) and (A.17), and because $W_{1}\leq W_{2}$, we obtain
$\frac{1}{N}\|G_{\mu}^{\mathrm{LOT}}-G_{\nu_{K}}^{\mathrm{LOT}}\|_{F}^{2}$ $\leq\frac{C_{N,\mathcal{X}}}{N}\sum\limits_{i=1}^{N}4\mathrm{diam}(\mathcal{X})LW_{1}(\mu^{(i)},\nu^{(i)}_{K})$
$\leq 4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\ \frac{1}{N}\sum\limits_{i=1}^{N}W_{2}(\mu^{(i)},\nu^{(i)}_{K})$
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{\left(\sum\limits_{i=1}^{N}W_{2}(\mu^{(i)},\nu^{(i)}_{K})\right)^{2}}$
$\leq 4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{N\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\nu^{(i)}_{K})}$
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\frac{1}{N}\sqrt{N\cdot N\varepsilon_{K}}$
$=4C_{N,\mathcal{X}}\mathrm{diam}(\mathcal{X})L\cdot\sqrt{\varepsilon_{K}}$
$=C_{N,\mathcal{X},L}\sqrt{\varepsilon_{K}},$
which concludes the proof. ∎
Proof of Proposition 4.7.
We start by following the classical guidelines in M-estimation in statistics by noticing that
$\mathcal{E}^{PCA}_{q}$ $=E(P_{K}^{\leq q})-E(P^{\leq q})$
$\leq E(P_{K}^{\leq q})+E_{K}(P^{\leq q})-E_{K}(P_{K}^{\leq q})-E(P^{\leq q})$ (A.18)
$\leq 2\sup\limits_{P\in\mathcal{P}_{q}}\bigl{|}E(P)-E_{K}(P)\bigr{|}$
where (A.18) is due to $E_{K}(P^{\leq q})-E_{K}(P_{K}^{\leq q})\geq 0$ as $P_{K}^{\leq q}$ is a minimizer for $E_{K}$.
Now, let us note $\mathbb{T}=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{T}^{(i)}}$ and $\mathbb{T}_{K}=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{T}^{(i)}_{K}}$. We have that:
$\Bigl{|}E(P)-E_{K}(P)\Bigr{|}$ $=\Bigl{|}-\langle\Sigma,P\rangle_{HS}-(-\langle\Sigma_{K},P\rangle_{HS})\Bigr{|}$
$=\Bigl{|}\frac{1}{N}\sum\limits_{i=1}^{N}\bigl{\langle}\overline{T}^{(i)}\otimes\overline{T}^{(i)},P\bigr{\rangle}_{HS}-\frac{1}{N}\sum\limits_{i=1}^{N}\langle\overline{T}^{(i)}_{K}\otimes\overline{T}^{(i)}_{K},P\bigr{\rangle}_{HS}\Bigr{|}$
$=\Bigl{|}\int\langle T\otimes T,P\rangle_{HS}\mathrm{d}\mathbb{T}(T)-\int\langle T\otimes T,P\rangle_{HS}\mathrm{d}\mathbb{T}_{K}(T)\Bigr{|}$
$=\Bigl{|}\int f_{P}(T)\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|},$
where $f_{P}:T\in L^{2}(\rho)\rightarrow\langle T\otimes T,P\rangle_{HS}\in\mathbb{R}$. In what follows, we aim to prove that this function $f_{P}$ is Lipschitz, that is
$|f_{P}(T_{1})-f_{P}(T_{2})|$ $=\Bigl{|}\langle T_{1}\otimes T_{1}-T_{2}\otimes T_{2},P\rangle_{HS}\Bigr{|}$
$=\Bigl{|}\langle T_{1}\otimes(T_{1}-T_{2})+(T_{1}-T_{2})\otimes T_{2},P\rangle_{HS}\Bigr{|}$
$\leq\|T_{1}\otimes(T_{1}-T_{2})+(T_{1}-T_{2})\otimes T_{2}\|_{HS}\cdot\|P\|_{HS}$
$\leq\bigl{(}\|T_{1}\|_{L^{2}(\rho)}\cdot\|T_{1}-T_{2}\|_{L^{2}(\rho)}+\|T_{2}\|_{L^{2}(\rho)}\cdot\|T_{1}-T_{2}\|_{L^{2}(\rho)}\bigr{)}\cdot\sqrt{q}$
$=\|T_{1}-T_{2}\|_{L^{2}(\rho)}\bigl{(}\|T_{1}\|_{L^{2}(\rho)}+\|T_{2}\|_{L^{2}(\rho)}\bigr{)}\cdot\sqrt{q}$
$\leq\|T_{1}-T_{2}\|_{L^{2}(\rho)}\cdot 2R\sqrt{q}$
where $R=\max\limits_{x\in\mathcal{X}}\|x\|$. The functional $\frac{f_{P}}{2R\sqrt{q}}$ is therefore 1-Lipschitz. We have that:
$\Bigl{|}E(P)-E_{K}(P)\Bigr{|}$ $=\Bigl{|}\int f_{P}(T)\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|}$
$=2R\sqrt{q}\Bigl{|}\int\frac{f_{p}(T)}{2R\sqrt{q}}\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|}$
$\leq 2R\sqrt{q}\Bigl{|}\sup\limits_{f\mbox{ is }1-\text{Lipschitz}}\int f(T)\mathrm{d}(\mathbb{T}-\mathbb{T}_{K})(T)\Bigr{|}$
$\leq 2R\sqrt{q}\cdot\mathcal{W}_{1}(\mathbb{T},\mathbb{T}_{K})$ (A.19)
$=2R\sqrt{q}\min\limits_{P~{}\text{coupling}}\sum\limits_{i,j=1}^{N}P_{ij}\|\overline{T}^{(i)}-\overline{T}^{(j)}_{K}\|_{L^{2}(\rho)}$
$\leq 2R\sqrt{q}\frac{1}{N}\sum\limits_{i=1}^{N}\|T^{(i)}-T^{(i)}_{K}\|_{L^{2}(\rho)}$
$\leq 2R\sqrt{q}\cdot\frac{1}{N}\sum\limits_{i=1}^{N}2\sqrt{\mathrm{diam}(\mathcal{X})L}W_{1}^{1/2}(\mu^{(i)},\nu_{K}^{(i)})$ (A.20)
$=4R\sqrt{q\mathrm{diam}(\mathcal{X})L}\cdot\frac{1}{N}\Biggl{(}\Bigl{(}\sum\limits_{i=1}^{N}W_{2}^{1/2}(\mu^{(i)},\nu_{K}^{(i)})\Bigr{)}^{4}\Biggr{)}^{1/4}$
$\leq 4R\sqrt{q\mathrm{diam}(\mathcal{X})L}\cdot\frac{1}{N}\Bigg{(}N^{3}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\nu_{K}^{(i)})\Biggr{)}^{1/4}$ (A.21)
$=4R\sqrt{q\mathrm{diam}(\mathcal{X})L}\cdot\varepsilon_{K}^{1/4}$
where (A.19) comes from the Kantorovich-Rubinstein formulation of $\mathcal{W}_{1}$ (see e.g. [51]). Inequality (A.20) follows from [17] and (A.21) follows from the inequality $\bigl{(}\sum\limits_{i=1}^{N}a_{i}\bigr{)}^{4}\leq N^{3}\sum\limits_{i=1}^{N}a_{i}^{4}$. We finally obtain that
$\mathcal{E}_{q}^{PCA}$ $\leq 2\sup\limits_{P\in\mathcal{P}_{q}}|E(P)-E_{K}(P)|$
$\leq 8R\sqrt{q\mathrm{diam}(\mathcal{X})L}\varepsilon_{K}^{1/4},$
which concludes the proof. ∎
Appendix B Link between the diagonalization of the covariance operator and the Gram matrix of inner-products
In this section, we show that diagonalizing the Gram matrix of inner-products is closely related to diagonalizing the covariance operator in a Hilbert space. Suppose we have elements $f_{1},\ldots,f_{N}$ belonging to a separable Hilbert space $\mathcal{H}$, endowed with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$. The covariance operator $\Sigma$ of the data is defined as:
$\forall h\in\mathcal{H},\qquad\Sigma(h)=\frac{1}{N}\sum\limits_{i=1}^{n}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}.$
We recall that $h\in\mathcal{H}$ is an eigenvector of $\Sigma$ associated to the eigenvalue $\lambda\in\mathbb{R}$ if
$\Sigma(h)=\frac{1}{N}\sum\limits_{i=1}^{n}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}=\lambda h$ (B.1)
Dividing by $\lambda$ in (B.1), one obtains:
$h=\sum\limits_{i=1}^{N}\frac{1}{N\lambda}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}=\sum\limits_{i=1}^{N}a_{i}f_{i},$ (B.2)
where $a_{i}=\frac{1}{N\lambda}\langle f_{i},h\rangle_{\mathcal{H}}\in\mathbb{R}$. Injecting (B.2) in (B.1) yields:
$\sum\limits_{i=1}^{N}f_{i}\langle f_{i},\sum\limits_{j=1}^{N}a_{j}f_{j}\rangle_{\mathcal{H}}=N\lambda\sum\limits_{i=1}^{N}a_{i}f_{i}$ (B.3)
Taking the inner-product of (B.3) with $f_{l}$ yields:
$\langle f_{l},\sum\limits_{i=1}^{N}f_{i}\langle f_{i},\sum\limits_{j=1}^{N}a_{j}f_{j}\rangle_{\mathcal{H}}\rangle_{\mathcal{H}}$ $=\langle f_{l},N\lambda\sum\limits_{i=1}^{N}a_{i}f_{i}\rangle_{\mathcal{H}}$
$\Leftrightarrow\sum\limits_{i=1}^{N}\langle f_{l},f_{i}\sum\limits_{j=1}^{N}a_{j}\langle f_{i},f_{j}\rangle_{\mathcal{H}}\rangle_{\mathcal{H}}$ $=N\lambda\sum\limits_{i=1}^{N}a_{i}\langle f_{l},f_{i}\rangle_{\mathcal{H}}$
$\Leftrightarrow\sum\limits_{i,j=1}^{N}a_{j}\langle f_{i},f_{j}\rangle_{\mathcal{H}}\langle f_{l},f_{i}\rangle_{\mathcal{H}}$ $=N\lambda\sum\limits_{i=1}^{N}a_{i}\langle f_{l},f_{i}\rangle_{\mathcal{H}}.$
If we note $K$ the $N\times N$ Gram matrix of inner-products, that is $K_{ij}=\langle f_{i},f_{j}\rangle_{\mathcal{H}}$, and $a=(a_{1},\cdots,a_{N})^{T}\in\mathbb{R}^{N}$, then we can rewrite the previous inequality with matrix notations:
$K^{2}a=N\lambda Ka$
and $a$ solves the following equation
$Ka=N\lambda a$
which can recovered by the diagonalization of the Gram matrix $K$. From $a$, one can recover the eigenelement $h$ of the covariance operator $\Sigma$ with (B.2).
Appendix C Additional numerical experiments
C.1 Synthetic dataset : an explicit formulation for the pairwise inner-products in the Gram matrix
Proposition C.1.
Let $\rho$ be an a.c. probability measure with compact support, defined for $X\sim\rho$ by $X=R\frac{Z}{\|Z\|}$, where $R\sim\mathrm{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,I_{d})$ are independent random variable. Let $(\mu^{(i)})_{i=1}^{N}$ be $N$ distributions defined by
$\mu^{(i)}=(\Sigma_{i}^{1/2}\mathrm{id}+b_{i})_{\#}\rho,$
where $\Sigma_{i}\in\mathbb{R}^{d\times d}$ is a positive semi-definite matrix and $b_{i}\in\mathbb{R}^{d}$.
(i) For the LOT embedding, we choose $\rho$ as the reference measure, for which $\phi(\mu^{(i)})=T_{\rho}^{\mu^{(i)}}-\mathrm{id}$. Then one has $\forall 1\leq i,j\leq N$,
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{L^{2}(\rho)}=\langle b_{i},b_{j}\rangle+\frac{1}{3d}\langle\Sigma_{i}^{1/2}-I_{d},\Sigma_{j}^{1/2}-I_{d}\rangle_{F}.$
(ii) For the KME, we consider the specific kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ given by $k(x,y)=x^{T}y+(x^{T}y)^{2}$, for which $\phi(\mu^{(i)})=\int k(x,\cdot)\mathrm{d}\mu^{(i)}(x)$. Then one has $\forall 1\leq i,j\leq N$,
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}=\langle b_{i},b_{j}\rangle+\langle b_{i},b_{j}\rangle^{2}+\frac{1}{3d}\langle\Sigma_{j},b_{i}b_{i}^{T}\rangle_{F}+\frac{1}{3d}\langle\Sigma_{i},b_{j}b_{j}^{T}\rangle_{F}+\frac{1}{9d^{2}}\langle\Sigma_{i},\Sigma_{j}\rangle_{F}.$
Proof of Proposition C.1.
First, we have $\mathbb{E}[X]=\mathbb{E}[R]\cdot\mathbb{E}\left[\frac{Z}{\|Z\|}\right]=\frac{1}{2}\cdot 0=0$. Then we also have that $\mathrm{Cov}[X]=\mathbb{E}[XX^{T}]=\mathbb{E}[R^{2}]\cdot\mathbb{E}\left[\frac{ZZ^{T}}{\|Z\|^{2}}\right]=\frac{1}{3}\cdot\frac{1}{d}I_{d}.$
(i) For LOT, we have
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{L^{2}(\rho)}$ $=\langle T_{\rho}^{\mu^{(i)}}-\mathrm{id},T_{\rho}^{\mu^{(j)}}-\mathrm{id}\rangle_{L^{2}(\rho)}$
$=\int\langle T_{\rho}^{\mu^{(i)}}(x)-x,T_{\rho}^{\mu^{(j)}}(x)-x\rangle\mathrm{d}\rho(x).$
Moreover, the optimal transport map between $\rho$ and $\mu^{(i)}$ is made explicit in (5.1). Note that it is optimal in the sense of (2.4) since $x\mapsto\Sigma_{i}^{1/2}x+b_{i}$ is the gradient of a convex function, and Brenier’s theorem [9] allows to conclude. Then
$\langle T_{\rho}^{\mu^{(i)}}-\mathrm{id},T_{\rho}^{\mu^{(j)}}-\mathrm{id}\rangle_{L^{2}(\rho)}$ $=\int\langle\Sigma_{i}^{1/2}x+b_{i}-x,\Sigma_{j}^{1/2}x+b_{j}-x\rangle\mathrm{d}\rho(x).$
Let us write $C_{i}=\Sigma_{i}^{1/2}-I_{d}$ for simplicity of notation, then since $\rho$ is centered:
$\langle T_{\rho}^{\mu^{(i)}}-\mathrm{id},T_{\rho}^{\mu^{(j)}}-\mathrm{id}\rangle_{L^{2}(\rho)}$ $=\int\langle C_{i}x+b_{i},C_{j}x+b_{j}\rangle\mathrm{d}\rho(x)$
$=\int(C_{i}x)^{T}C_{j}x\mathrm{d}\rho(x)+\int b_{i}^{T}b_{j}\mathrm{d}\rho(x)$
$\quad+\int(C_{i}x)^{T}b_{j}\mathrm{d}\rho(x)+\int(C_{j}x)^{T}b_{i}\mathrm{d}\rho(x)$
$=b_{i}^{T}b_{j}+\int x^{T}C_{i}C_{j}x\mathrm{d}\rho(x).$
Finally, using that $\mathbb{E}[XX^{T}]=\frac{1}{3d}I_{d}$,
$\int x^{T}C_{i}C_{j}x\mathrm{d}\rho(x)$ $=\int\mathrm{Tr}\bigl{(}x^{T}C_{i}C_{j}x\bigr{)}\mathrm{d}\rho(x)$
$=\mathrm{Tr}\Bigl{(}\int C_{i}C_{j}xx^{T}\mathrm{d}\rho(x)\Bigr{)}$
$=\mathrm{Tr}\Bigl{(}C_{i}C_{j}\int xx^{T}\mathrm{d}\rho(x)\Bigr{)}$
$=\mathrm{Tr}\left(C_{i}C_{j}\frac{1}{3d}I_{d}\right)$
$=\frac{1}{3d}\mathrm{Tr}\bigl{(}C_{i}C_{j}\bigr{)}.$
And we get
$\langle T_{\rho}^{\mu^{(i)}}-\mathrm{id},T_{\rho}^{\mu^{(j)}}-\mathrm{id}\rangle_{L^{2}(\rho)}=\langle b_{i},b_{j}\rangle+\frac{1}{3d}\langle\Sigma_{i}^{1/2}-I_{2},\Sigma_{j}^{1/2}-I_{2}\rangle_{F}.$
(ii) For the KME embedding,
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}$ $=\left\langle\int k(x,\cdot)\mathrm{d}\mu^{(i)}(x),\int k(y,\cdot)\mathrm{d}\mu^{(j)}(y)\right\rangle_{\mathcal{H}}$
$=\iint\langle k(x,\cdot),k(y,\cdot)\rangle_{\mathcal{H}}\ \mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=\iint k(x,y)\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=\iint\left[x^{T}y+(x^{T}y)^{2}\right]\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=\iint\left[x^{T}y+x^{T}yy^{T}x\right]\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=b_{i}^{T}b_{j}+\iint x^{T}yy^{T}x\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=b_{i}^{T}b_{j}+\iint\mathrm{Tr}\bigl{(}x^{T}yy^{T}x\bigr{)}\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=b_{i}^{T}b_{j}+\iint\mathrm{Tr}\bigl{(}yy^{T}xx^{T}\bigr{)}\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)$
$=b_{i}^{T}b_{j}+\mathrm{Tr}\bigl{(}\iint yy^{T}xx^{T}\mathrm{d}\mu^{(i)}(x)\mathrm{d}\mu^{(j)}(y)\bigr{)}$
$=b_{i}^{T}b_{j}+\mathrm{Tr}\Bigl{(}\int yy^{T}\mathrm{d}\mu^{(j)}(y)\int xx^{T}\mathrm{d}\mu^{(i)}(x)\Bigr{)}$
Now, since $\mu^{(i)}=(\Sigma_{i}^{1/2}\mathrm{id}+b_{i})_{\#}\rho$, we have that :
$\int xx^{T}\mathrm{d}\mu^{(i)}(x)$ $=\int\bigl{(}\Sigma_{i}^{1/2}x+b_{i}\bigr{)}\bigl{(}\Sigma_{i}^{(1/2}x+b_{i}\bigr{)}^{T}\mathrm{d}\rho(x)$
$=\int\bigl{(}\Sigma_{i}^{1/2}xx^{T}\Sigma_{i}^{1/2}+\Sigma_{i}^{1/2}xb_{i}^{T}+b_{i}x^{T}\Sigma_{i}^{1/2}+b_{i}b_{i}^{T}\bigr{)}\mathrm{d}\rho(x)$
$=\Sigma_{i}^{1/2}\int xx^{T}\mathrm{d}\rho(x)\Sigma_{i}^{1/2}+b_{i}b_{i}^{T}$
$=\frac{1}{3d}\Sigma_{i}+b_{i}b_{i}^{T}.$
This gives :
$\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}}$ $=b_{i}^{T}b_{j}+\mathrm{Tr}\Bigl{(}\bigl{(}\frac{1}{3d}\Sigma_{i}+b_{i}b_{i}^{T}\bigr{)}\bigl{(}\frac{1}{3d}\Sigma_{j}+b_{j}b_{j}^{T}\bigr{)}\Bigr{)}$
$=b_{i}^{T}b_{j}+\mathrm{Tr}\Bigl{(}\frac{1}{9d^{2}}\Sigma_{i}\Sigma_{j}+\frac{1}{3d}\Sigma_{i}b_{j}b_{j}^{T}+\frac{1}{3d}\Sigma_{j}b_{i}b_{i}^{T}+b_{i}b_{i}^{T}b_{j}b_{j}^{T}\Bigr{)}$
$=\langle b_{i},b_{j}\rangle+\langle b_{i},b_{j}\rangle^{2}+\frac{1}{3d}\langle\Sigma_{j},b_{i}b_{i}^{T}\rangle_{F}+\frac{1}{3d}\langle\Sigma_{i},b_{j}b_{j}^{T}\rangle_{F}+\frac{1}{9d^{2}}\langle\Sigma_{i},\Sigma_{j}\rangle_{F}$
Figure 5 depicts the first components of PCA after KME on both quantization steps. As for LOT embedding, we see that for $K$ as small as 32, the PCA visualizations with the two quantization methods look highly similar to the true PCA.
Figure 5: Synthetic dataset on shifts and scalings. Projection of the data onto the first two components of PCA after $\overline{K}$-KME (top) and $\tilde{K}$-KME (bottom) and comparison to the KME PCA (right) computed from the true Gram matrix (see Prop.C.1.
C.2 Visualization of the flow cytometry dataset
In this section, we show in Figures 6 and 7 the projections of the embedded data (by either $K$-LOT or $K$-KME) on the first components of PCA. We see that the first two components already allow to discriminate the flow cytometry measurements according to their labels.
Figure 6: Flow cytometry dataset. Projection of the $N=108$ measures on the first components of PCA after embedding with $K$-LOT.
Figure 7: Flow cytometry dataset. Projection of the $N=108$ measures on the first components of PCA after embedding with $K$-KME with RBF kernel with $\sigma=1$.
C.3 Performance of the methods for different values of $K$
In this section, we display additional experiments on the flow cytometry and Airbus datasets, completing Tables 1 and 2 for different values of $K$.
Table 3: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=16$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 85 23
$\tilde{K}$-LOT 100 81 103
Random subset of size $K$ + LOT 100 77 22
Random subset of size $K$ + KME 100 67 14
$\overline{K}$-KME 100 83 15
$\tilde{K}$-KME 100 69 96
KME with RFF 73 44 4524
$K$-Nys-KME 100 71 12
Table 4: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=32$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 94 25
$\tilde{K}$-LOT 100 81 166
Random subset of size $K$ + LOT 100 77 23
Random subset of size $K$ + KME 100 69 32
$\overline{K}$-KME 100 83 34
$\tilde{K}$-KME 100 69 174
KME with RFF 75 44 4701
$K$-Nys-KME 100 73 19
Table 5: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=128$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 88 32
$\tilde{K}$-LOT 100 81 555
Random subset of size $K$ + LOT 100 79 28
Random subset of size $K$ + KME 100 73 376
$\overline{K}$-KME 100 77 387
$\tilde{K}$-KME 100 71 909
KME with RFF 92 44 5676
$K$-Nys-KME 100 73 65
Table 6: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=16$.
Method Accuracy Time (s)
$\overline{K}$-LOT 88 15
$\tilde{K}$-LOT 88 228
Random subset of size $K$ + LOT 67 21
Random subset of size $K$ + KME 69 523
$\overline{K}$-KME 76 219
$\tilde{K}$-KME 68 732
$K$-Nys-KME 65 169
Table 7: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=32$.
Method Accuracy Time (s)
$\overline{K}$-LOT 89 17
$\tilde{K}$-LOT 89 249
Random subset of size $K$ + LOT 72 28
Random subset of size $K$ + KME 70 2031
$\overline{K}$-KME 67 2792
$\tilde{K}$-KME 67 2247
$K$-Nys-KME 65 172
Table 8: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=128$.
Method Accuracy Time (s)
$\overline{K}$-LOT 88 32
$\tilde{K}$-LOT 88 390
Random subset of size $K$ + LOT 80 57
Random subset of size $K$ + KME 70 31915
$\overline{K}$-KME 67 31765
$\tilde{K}$-KME 66 32287
$K$-Nys-KME 65 437
C.4 Earth image dataset
We display in Figure 8 a few samples of the earth image dataset.
(a) With wind turbine
(b) Without wind turbines
Figure 8: Earth image dataset. Examples of images sampled from the Airbus dataset.